### v1

# LLM-Based Compact Reranking with Document Features for Scientific Retrieval

Abstract
Scientific retrieval is essential for advancing academic discovery. Within this process, document reranking plays a critical role by refining first-stage retrieval results. However, standard LLM listwise reranking faces unique challenges in the scientific domain. First-stage retrieval is often suboptimal in the scientific domain, so relevant documents are ranked lower. Moreover, conventional listwise reranking inputs the full text of candidates into the context window, limiting the number of candidates that can be considered. As a result, many relevant documents are excluded before reranking, constraining overall retrieval performance. To address these challenges, we explore compact document representations based on semantic features (e.g., categories, sections and keywords) and propose CoRank, a training-free, model-agnostic reranking framework for scientific retrieval. It presents a three-stage solution: (i) offline extraction of document-level features, (ii) coarse reranking using these compact representations, and (iii) fine-grained reranking on full texts of the top candidates from (ii). This hybrid design provides a high-level abstraction of document semantics, expands candidate coverage, and retains critical details required for precise ranking. Experiments on LitSearch and CSFCube show that CoRank significantly improves reranking performance across different LLM backbones (nDCG@10 from $32.0$ to $39.7$). Overall, these results highlight the value of information extraction for reranking in scientific retrieval. Code will be publicly available.
LLM-Based Compact Reranking with Document Features for Scientific Retrieval
Figure 1: Instead of directly reranking full documents, we extract features, rerank a larger candidate pool with them, and finally refine the shortlist with full documents.
1 Introduction
Scientific retrieval (Lawrence et al., 1999; White et al., 2009) is crucial for scientific discovery. While current retrievers are effective at retrieving broadly relevant scientific papers, they often struggle to differentiate between documents covering similar topics (Sciavolino et al., 2021; Liu et al., 2021), making fine-grained relevance estimation essential. Therefore, the reranking stage (Carbonell and Goldstein-Stewart, 1998; Kurland and Lee, 2005) is particularly important (Gao et al., 2021b), as it refines the first-stage retrieval to better distinguish between closely ranked documents.
Recently, large language models (LLMs) (Grattafiori et al., 2024; OpenAI et al., 2024) have significantly advanced document reranking, particularly through their application to listwise reranking (Sun et al., 2023a; Ma et al., 2023). In this setting, LLMs jointly assess a set of retrieved candidates within their context window and generate a reordered ranking based on their relevance. Previous works (Pradeep et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c; Ren et al., 2024) show that LLM-based listwise rerankers outperform prior embedding-based approaches (Nogueira et al., 2019b, 2020) on benchmarks like BEIR (Thakur et al., 2021).
The standard practice in LLM listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) is to input the full text of each candidate document into the context window. However, this approach faces key limitations in scientific retrieval. In the scientific domain, retrievers often show limited performance (Kim et al., 2023; Kang et al., 2024a, b), so truly relevant documents may not rank high enough in the first-stage retrieval. Moreover, due to the substantial token overhead of full text representation and finite context length, rerankers can only operate on a limited number (typically 20 per prompt) (Sun et al., 2023a; Ma et al., 2023) of candidates. As a result, when first-stage retrieval is suboptimal, the reranking performance is inherently constrained (Reddy et al., 2023).
To address these limitations, we explore an alternative document representation that is both compact and informative. Instead of using full text, we investigate an information extraction (IE)–based approach (Niklaus et al., 2018; Zhou et al., 2022; Liu et al., 2022), representing each scientific paper using high-level features such as categories, sections, and keywords. We find that this representation significantly improves per-document token efficiency, enabling a larger pool of candidates to fit within the LLM’s context window. This, in turn, makes the reranking process more robust to suboptimal first-stage scientific retrieval results. It also simplifies the input by filtering out irrelevant details, making it easier for LLMs to interpret. However, since IE cannot guarantee complete coverage of all relevant content, full-text inputs still serve as a valuable complement when fine-grained relevance comparisons are required.
Motivated by these insights, we propose CoRank, a training-free, model-agnostic reranking framework for science retrieval. CoRank consists of three stages: (i) Offline Preprocessing: extract high-level semantic features like categories and keywords from unstructured scientific documents; (ii) Coarse Reranking: use these compact representations for an initial ranking and select a subset of top candidates; (iii) Fine-grained Reranking: rerank the top candidates using full scientific documents to recover the potentially missing details during information extraction. Our design enhances the robustness against suboptimal first-stage retrieval with compact feature representation and maximizes the overall effectiveness with final full text reranking.
We evaluate CoRank on two scientific retrieval benchmarks: LitSearch (Ajith et al., 2024) and CSFCube (Mysore et al., 2021). Empirical experiments show that across various LLM backbones, CoRank achieves an absolute improvement of $+11.5$ nDCG@10 without the sliding window strategy, and retains a $+3.9$ gain with it. These results demonstrate significant effectiveness of our method for reranking in scientific retrieval. Overall, our contributions are threefold:
  1. # 1
    We are the first to reveal the unique limitations of LLM listwise reranking in scientific retrieval.
  2. # 2
    We are the first to explore semantic features as compact representations in reranking and propose CoRank, a framework based on this design.
  3. # 3
    We show that CoRank significantly improves reranking performance in scientific retrieval.
These findings underscore the value of compact feature extraction as a pre-analysis step for improving the reranking performance in scientific retrieval.
Figure 2: Overview of our feature extraction pipeline: from unstructured documents, we apply zero-shot LLM information extraction to obtain document features including categories, sections, pseudo queries, and keywords, which are then combined into compact representations.
2 Preliminary Analysis
In this section, we first define listwise reranking, then show the limitations of current full text-based methods and explore compact and informative document representation options for reranking in the scientific domain.
2.1 Problem Definition
Reranking Input.
Given a query $q$ and a large corpus $\mathcal{P}_{n}$ of $n$ scientific papers, a first-stage retriever selects a ranked list of $m$ candidate documents $\mathcal{C}=\{p_{1},p_{2},\dots,p_{m}\}\subset\mathcal{P}_{n}$. Typically we have $m\ll n$ due to the scale of the corpus and the finite capacity of rerankers. This candidate list $\mathcal{C}$ is then passed to a reranking model, which aims to reorder the documents such that more relevant ones are ranked higher.
Reranking Objective.
The goal of document reranking is to find a permutation $\sigma:\{1,\dots,m\}\!\to\!\{1,\dots,m\}$ whose application
$\mathcal{C}^{\prime}\;=\;[\,p_{\sigma(1)},\,p_{\sigma(2)},\dots,p_{\sigma(m)}\,]$
orders documents in more accurate descending relevance to $q$. The quality is often (Sun et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c) evaluated using top-$k$ metrics such as nDCG@10, with small $k$ values that stress accuracy at the top of the ranked list.
LLM Listwise Reranking.
Large language models can inspect $m$ candidate documents in their context window and generate a permutation that reflects their relevance ordering. Formally, the model acts as follows:
$\sigma_{\text{LLM}}\;=\;\mathrm{LLM}(q,\,\mathcal{C})\;\in\;S_{m},$
where $S_{m}$ is the set of all permutations on $\{1,\dots,m\}$. Applying this permutation to the candidate list yields
$\mathcal{C}^{\prime}\;=\;\bigl{[}p_{\sigma_{\text{LLM}}(1)},\;p_{\sigma_{\text{LLM}}(2)},\;\dots,\;p_{\sigma_{\text{LLM}}(m)}\bigr{]}.$
2.2 Current Limitations
The standard approach in LLM-based listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) feeds the full text of each candidate into the model’s context window. While this allows LLMs to capture the complete content of each document, it incurs some limitations, especially in scientific retrieval.
Suboptimal First-Stage Retrieval.
Unlike in general-domain scenarios, first-stage retrievers—whether sparse (Robertson et al., 2009; Nogueira et al., 2019a) or dense (Izacard et al., 2021; Wang et al., 2022)—struggle to generalize to the scientific domain (Thakur et al., 2021; Bonifacio et al., 2022). This is due to the complexity of long-tail concepts (Kang et al., 2024b, 2025) and the lack of large-scale supervised training data (Bonifacio et al., 2022; Li et al., 2023). Consequently, their retrieval performance is limited, and truly relevant documents are often ranked much lower (Ajith et al., 2024).
Token Consumption.
Full text representation also introduces significant token overhead, with individual scientific papers often consuming hundreds or even thousands of tokens (Thakur et al., 2021; Ajith et al., 2024; Mysore et al., 2021). At the same time, LLMs have a limited effective context length (Dai et al., 2019; Xiong et al., 2024) and are known to suffer from issues such as positional bias for long inputs (Liu et al., 2024a; Tian et al., 2024). As a result, full-text rerankers are constrained to operate on a small number of candidates, typically 20 documents per input prompt (Sun et al., 2023a; Ma et al., 2023; Pradeep et al., 2023a, b).
When first-stage retrieval is suboptimal and reranking operates over a narrow candidate set, the overall reranking performance is significantly constrained.
2.3 Semantic Features as Alternatives
Given the limitations of full-text representations in scientific-domain reranking, we explore an alternative approach, which is based on document-level semantic features such as categories and keywords. The intuition is that these features are both significantly more concise and capable of preserving the core semantics. Therefore, more candidate documents can be effectively considered during reranking. Specifically, we examine the following four types of document semantic features:
Category.
Categories (Sun et al., 2023b; Zhang et al., 2024) offer a high-level topical overview of scientific documents. Each document is assigned a three-level hierarchical category path in the format {Category} → {Subcategory} → {Subsubcategory}, which provides a broad-to-specific classification on its topic.
Section.
Sections (Zhou et al., 2023) consist of multiple subtitle-style strings, each summarizing a major part of the scientific document. They capture the internal structure of the document and serve as mid-level semantic signals that enhance category-level summaries.
Keyword.
Keywords (Rose et al., 2010; Lee et al., 2023) are terms or entities that represent fine-grained lexical concepts within a document. They provide the most specific information among different granularities.
Pseudo Query.
Pseudo queries (Sachan et al., 2022; Kang et al., 2025) are synthetic user questions based on the content. This feature offers a unique query-aligned perspective of the document, simulating how the document might be retrieved in real-world scenarios.
These features cover most common types of document-level IE and vary in granularity and style. All semantic features are extracted using LLM-based zero-shot information extraction. The prompt templates used for this extraction are provided in Appendix A.
Among these features, sections, keywords, and pseudo queries have multiple elements for each document. To reduce noise and focus on the most relevant content, we apply an adaptive selection strategy at inference time. We compute dense embedding similarities between the query and each extracted element and retain only the most relevant: 5 keywords (from 30), 1 pseudo query (from 20), and 1 section (from 3). This improves content relevance while minimizing token overhead. Ablation of this strategy can be found in Section 4.2.
Figure 3: Token efficiency and performance comparison across different document representations. (a) Per-document token lengths distribution. (b) nDCG@10 scores for different number of documents in the context window of single LLM input (c) Context token overhead to reach equal reranking performance.
However, relying on a single feature often lacks representational power. For example, using only the category tends to be too coarse-grained and offers limited discriminative ability, while keywords alone may lack sufficient context or background information. To address this, we represent each document using combinations of features. Specifically, we explore the following four configurations of different overall granularities.
  1. Form 1.
    Pseudo Query
  2. Form 2.
    Category
  3. Form 3.
    Category + Section
  4. Form 4.
    Category + Section + Keywords
Ablation studies on the effect of each specific component can further be found in Section 4.2. The overview of representation construction is presented in Figure 2.
2.4 Empirical Validation
To evaluate the effectiveness of our feature-based representations, we conduct a series of experiments on the LitSearch (Ajith et al., 2024), using GPT-4.1-mini (OpenAI, 2024) as the reranking backbone. Semantic features are extracted using Qwen3-8B-Instruct (Qwen et al., 2024). Following Gangi Reddy et al. (2024), we use Contriever (Izacard et al., 2021) as the first-stage retriever and the similarity encoder for adaptive selection.
Token Length Per Document.
We first evaluate the token cost of different representations. As shown in Figure 3 (a), full-text inputs consume around 200 tokens per document on average, with some exceeding 5,000 tokens, limiting the number of documents that can fit into the context window. In contrast, semantic feature-based representations (Forms 1-Form 4) are much more compact, each averaging between 10 and 50 tokens.
Number of Documents in a Single Input.
We then evaluate how different representations perform as the number of candidate documents increases (up to 200) within a 32k-token context window. As shown in Figure 3 (b), all representations improve as more documents are included. In the range of 0 to 80 candidates, full-text representations perform better than feature-based ones. This is expected, since full text retains complete document information, though at a much higher token cost. However, beyond 80 candidates, full-text inputs frequently exceed the context length limit. In contrast, feature-based methods, especially Form 4, remain compact and can handle more candidates without issue, while still achieving comparable performance.
Token Cost for Equal Reranking Performance.
Finally, we examine how many context tokens are needed to achieve the same reranking performance across different representations. Using results from earlier experiments, we estimate the total token cost by combining the number of documents required with the average token count per document. We take full-text reranking with 20 documents as the 100% performance base, and compare how many tokens each method needs to reach 70%, 85%, 100%, 115%, and 130% of that level. As shown in Figure 3 (c), compact feature-based representations can match full-text performance with significantly fewer tokens. For instance, reaching 100% performance requires around 3,500 tokens for full text, but only 1,000 or even a few hundred tokens for feature-based inputs.
2.5 Discussion
The results above demonstrate that feature-based document representations offer clear advantages over full-text inputs in reranking. They can express the core content of a document using far fewer tokens, allowing significantly more retrieved candidates to be included within the same token overhead. This is especially valuable for recovering relevant documents that were assigned lower scores by the first-stage retriever. Considering both token efficiency and reranking performance, we select Form 4 as our feature-based representation.
However, in the second set of experiments where the number of documents is held constant, we also observe that feature-based representations may underperform full text. After all, offline information extraction may omit subtle but important cues present in the full text.
Therefore, these results motivate a hybrid reranking strategy: using semantic features in the early stage to cover a broader range of candidates, followed by a refinement stage that reranks the top candidates using full-text inputs for more in-depth comparison.
3 Methodology
Based on the preliminary analysis, we propose a hybrid reranking framework in scientific domain. It proceeds in three steps: (i) offline extraction of structured semantic features; (ii) coarse-grained reranking using the features to select a subset of top candidates; and (iii) fine-grained reranking over the subset with full text inputs.
3.1 Compact and Informative Representation
To construct compact, feature-based representations for each scientific document, we first perform document-level information extraction offline. As described in Section 2.3, we use zero-shot information extraction to obtain target semantic features: category, section, and keyword. Formally, for a given document $p_{i}$, we obtain:
$\text{LLM}(p_{i})=\text{Category}_{i},\,[\text{Section}_{i}^{j}],\,[\text{Keyword}_{i}^{j}]$
These features are designed to capture the core semantics of scientific documents in a token-efficient manner. The extraction is performed offline, and the results are cached and reused at inference time to avoid any additional runtime delay. The prompt templates for extracting different features can be found in Appendix A.
3.2 Coarse Reranking w. Compact Features
Given a query $q$, we obtain a document relevance ranking $\mathcal{C}=[p_{1},\,p_{2},\,\dots,\,p_{m}]$ from a first-stage retriever. We then replace each document with a compact semantic representation and apply LLM-based listwise reranking to identify a high-quality subset.
For sections and keywords, we apply adaptive selection to select only those most relevant to the query. Specifically, for each unit (single section or keyword) $u$, we compute the cosine similarity with text embedding and select the top-5 keywords and the most relevant section. (Analysis on the number of keyword used can be found in Section 4.4) The final feature-based representation $r_{i}$ is formed as:
$r_{i}=\text{Concat}\left[\text{Category}_{i},\,\text{Section}_{i}^{*},\,\text{Keywords}_{i}^{*}\right]$
where indicates adaptively selected elements.
With the feature-based representations $\mathcal{R}=[r_{1},\,r_{2},\,\dots,\,r_{m}]$, we then use the LLM to perform listwise reranking, producing a permutation $\sigma_{\text{feat}}=\text{LLM}(q,\,\mathcal{R})\in S_{m}$. Applying this permutation to the original candidate list $\mathcal{C}=[p_{1},\,p_{2},\,\dots,\,p_{m}]$ yields the reranked output for the coarse reranking:
$\mathcal{C}_{\text{feat}}=[p_{\sigma_{\text{feat}}(1)},\,p_{\sigma_{\text{feat}}(2)},\,\dots,\,p_{\sigma_{\text{feat}}(m)}]$
We then keep the top-$k$ documents from this list to form the seed set for full text reranking:
$\mathcal{C}_{\text{seed}}=[p_{\sigma_{\text{feat}}(1)},\,\dots,\,p_{\sigma_{\text{feat}}(k)}],\quad k<m$
Coarse reranking with compact document representations greatly expands the number of candidates that can be considered in the same LLM input. This broader coverage helps recover scientific documents that were initially assigned low scores by the first-stage retriever.
3.3 Fine-grained Reranking w. Full Text
In the second reranking stage, we refine the ranking over the seed set of candidates using full documents. Specifically, we take the seed set $\mathcal{C}_{\text{seed}}=[p_{1}^{\prime},\,p_{2}^{\prime},\,\dots,\,p_{k}^{\prime}]$, (we use to distinguish them from the first-stage input) obtained from the previous stage, and replace each compact representation with its original document text.
Let $\mathcal{T}=[t_{1},\,t_{2},\,\dots,\,t_{k}]$ denote the full text of the selected documents, where $t_{i}$ corresponds to the full text content of passage $p_{i}^{\prime}$. We then once again use LLM to perform listwise reranking:
$\sigma_{\text{text}}=\text{LLM}(q,\,\mathcal{T})\in S_{k}$
With the permutation $\sigma_{\text{text}}$ we get the final ranking:
$\mathcal{C}_{\text{final}}=[p_{\sigma_{\text{text}}(1)}^{\prime},\,\dots,\,p_{\sigma_{\text{text}}(k)}^{\prime}]$
Fine-grained reranking with full-text inputs recovers details that may be lost during the information extraction process. Since the candidate set has already been narrowed down, full documents can now be used without exceeding the LLM’s context limit.
Overall, this hybrid strategy effectively addresses the challenges of LLM-based listwise reranking in the scientific domain. The coarse reranking stage expands the reranking pool, improving robustness to suboptimal first-stage retrieval in the scientific domain. The fine-grained reranking stage, in turn, preserves sufficient detail for precise relevance comparisons.
Model Strategy LitSearch CSFCube
N@5 N@10 M@5 M@10 R@5 R@10 N@5 N@10 M@5 M@10 R@5 R@10
Initial Retriever
Contriever Dense Retrieval 12.6 14.2 11.0 11.7 16.4 21.2 22.0 23.6 5.8 8.6 9.5 16.4
Supervised, Model-Specific Methods
[…]
Zero-Shot, Model-Agnostic Reranking Strategies
[…]
Table 1: Reranking performance on the LitSearch and CSFCube reflected by nDCG (N@k), MAP (M@k) and Recall (R@k). We compare CoRank with supervised reranking models and zero-shot, model-agnostic strategy.
4 Experiments
In this section, we first outline the experimental setup, then present the reranking results, followed by ablation studies and further analysis. Detailed qualitative studies can be found in Appendix B.
4.1 Experimental Setup
Datasets & Metric.
We evaluate different reranking methods on two high-quality scientific retrieval benchmarks: LitSearch (Ajith et al., 2024) and CSFCube (Mysore et al., 2021). The former is a benchmark of expert-annotated complex literature queries targeting recent ML and NLP papers. The latter is a human-annotated testbed for faceted query-by-example retrieval. Since reranking focuses on the top results, we follow standard practice (Sun et al., 2023a; Ma et al., 2023; Gangi Reddy et al., 2024; Liu et al., 2024c), reporting metrics for @5 and @10. We include nDCG, MAP, and Recall as our evaluation metrics.
Models & Hyperparameters.
For the reranking backbones, we use Qwen3-32B-Instruct (Qwen et al., 2024) as the representative open-source model, Gemini 2.0 Flash (Google DeepMind, 2025) and GPT-4.1-mini (OpenAI, 2024) as proprietary examples. For generative parameters, we use a temperature of 1.0 and a fixed random seed of 42 for reproducibility. The API costs are detailed in Appendix C.
For semantic feature extraction, we find that small open-source models are sufficiently effective; we therefore adopt Qwen3-8B-Instruct (Qwen et al., 2024) for all extraction tasks (Templates in Appendix A). For both first-stage retrieval and semantic similarity scoring for filtering features, we use Contriever (Izacard et al., 2021) following Gangi Reddy et al. (2024).
Baselines.
For supervised, model-specific baselines, we compare against RankVicuna (Pradeep et al., 2023a), RankZephyr (Pradeep et al., 2023b), and RankMistral (Liu et al., 2024c). These models are trained on large-scale general-domain datasets such as MS MARCO (Bajaj et al., 2016) and RankGPT (Sun et al., 2023a) to enhance reranking performance.
For model-specific baselines, we use vanilla listwise reranking (Sun et al., 2023a) with zero-shot instruction-following. We also consider the sliding window strategy (Sun et al., 2023a; Ma et al., 2023), a test-time scaling (Xia et al., 2025) technique designed to enhance reranking performance. We report results both with and without the use of sliding windows. Although CoRank is orthogonal to this technique, we include a comparison in Section 4.3 to further show our effectiveness.
Reranking Parameters.
We follow the standardized setup (Sun et al., 2023a; Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) from prior work to unify reranking parameters. For vanilla reranking without the sliding window strategy, we include 20 full-text documents within the context. An exception is RankMistral (Liu et al., 2024c), which benefits from long-context training and is able to process 100 full-text documents in a single input. When using the sliding window strategy, we rank a total of 100 full-text documents by applying a window size of 20 with a step size of 10. For CoRank, we include 200 compact representations in the coarse reranking stage, followed by 20 full-text documents in the fine-grained reranking stage.
4.2 Main Results
Reranking Performance.
We begin by evaluating the performance of different reranking methods on the target academic retrieval benchmarks, as shown in Table 1.
First, we observe that all reranking methods outperform the first-stage dense retriever baseline, confirming the critical role of reranking in scientific document retrieval. However, while model-specific rerankers such as RankVicuna (Pradeep et al., 2023a) and zero-shot listwise baselines do yield noticeable improvements, their gains are relatively modest compared to results in general-domain settings (Sun et al., 2023a), particularly when the sliding window strategy is not used.
In contrast, our proposed method, CoRank, is fully zero-shot, model-agnostic, and training-free, yet achieves strong and consistent gains across different LLM backbones. Notably, without using sliding windows, CoRank improves average nDCG@10 from 27.3 to 38.8, a relative improvement of over 40%. Even when combined with sliding window inputs, it still yields an average gain of +3.9 nDCG@10.
We also find that both CoRank and the sliding window strategy independently lead to substantial performance gains (ours being larger and more efficient; see Section 4.3). This is likely because both methods expand the reranking scope, which is particularly valuable in scientific domains considering the suboptimal first-stage retrieval. These findings reinforce our earlier analysis of current limitations presented in Section 2.2.
Ablation Studies.
To evaluate the contribution of each component in our design, we conduct ablation studies on the LitSearch (Ajith et al., 2024) dataset using GPT-4.1-mini (OpenAI, 2024) as the backbone.
Specifically, we remove individual semantic features—category, section, and keywords—from the compact representation to assess their relative importance. In addition, we ablate two key components of our pipeline: adaptive selection and fine-grained reranking, to understand their impact on overall performance.
Model Strategy N@10 M@10 R@10
Vanilla Full Text 26.1 25.2 27.3
CoRank – Category 43.4 41.9 45.9
CoRank – Section 45.4 43.8 48.2
CoRank – Keywords 44.6 42.8 48.0
CoRank – Selection 45.5 44.1 48.2
CoRank – Fine. Rank. 41.1 38.7 46.7
CoRank Full 46.3 44.7 49.4
Table 2: Ablation results showing the impact of removing individual features and two key design components.
As shown in Table 2, removing any single semantic component (category, section, or keywords) consistently degrades reranking performance, confirming that each feature contributes complementary information. Category captures coarse-grained background context that tends to align more easily with the query but offers less specific detail. In contrast, section and keywords reflect finer-grained semantics that are harder to match but, when relevant, provide highly targeted signals. This balance underscores the value of combining semantic features across different levels of abstraction.
Additionally, we find that both adaptive selection and fine-grained reranking contribute to the final performance. Adaptive selection ensures that the most query-relevant features are retained, while the fine-grained reranking stage substantially compensates for potential information loss in compact representations by recovering more nuanced evidence from the full text.
4.3 Comparison with Sliding Window
Sliding window strategy (Sun et al., 2023a) is a widely used (Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) method for overcoming the context length limitations of LLMs in listwise reranking. Instead of ranking all candidates in a single prompt, it partitions the candidate list into overlapping windows and reranks each window independently from the bottom up. This method has been shown to notably improve reranking performance (Sun et al., 2023a).
As noted in Section 4.1, we evaluate both the standalone and combined use of sliding windows with our method. While CoRank and sliding window are orthogonal and compatible, isolating them in comparison allows us to clearly demonstrate the superior efficiency and effectiveness of our approach when used alone.
Method N@10 M@10 R@10 Token Usage Cost
Vanilla 26.0 25.0 27.3 1.01M \$0.40
Sliding 40.8 39.1 44.2 9.06M \$3.62
CoRank 42.2 40.6 45.2 3.60M \$1.44
Both 43.8 41.8 48.1 11.65M \$4.66
Table 3: Comparison of performance, token usage, and API cost between CoRank and Sliding Window.
We evaluate CoRank and the sliding window strategy on LitSearch and CSFCube on three tested models, reporting nDCG@10, total token consumption, and API cost for GPT-4.1-mini. As shown in Table 3, CoRank not only achieves better average reranking quality (nDCG@10 improved by 1.4), but also requires only 40% of the token budget compared to sliding windows. These results demonstrate that CoRank provides an efficient yet effective alternative for expanding reranking range. Moreover, combining these two methods can further boost performance, offering an advanced solution for complex retrieval scenarios.
4.4 Hyperparameter Study
We also study the effect of key hyperparameters in CoRank, using the GPT-4.1-mini (OpenAI, 2024) reranker on the LitSearch (Ajith et al., 2024) dataset. We report results in terms of nDCG@10, MAP@10, Recall@10, token usage, and estimated API cost.
Number of Keywords.
With its default configuration, CoRank extracts 30 keywords from each document and selects the top 5 most relevant ones based on cosine similarity from text embedding. These selected keywords are then concatenated into the document representation. Here, we evaluate the impact of the keyword selection on coarse reranking performance by varying the number of concatenated keywords: 0, 1, 3, 5, 10, 15, and 20. The results are in Table 4.
# Keyword N@10 M@10 R@10 Token Usage Cost
0 keyword 38.9 36.5 44.2 1.80M \$0.72
1 keyword 40.4 38.0 46.3 1.95M \$0.78
3 keywords 40.1 37.9 44.9 2.26M \$0.90
5 keywords 41.1 38.7 46.7 2.57M \$1.03
10 keywords 42.4 40.4 46.5 3.33M \$1.33
15 keywords 42.6 40.1 48.6 4.10M \$1.64
20 keywords 42.4 40.2 47.4 4.87M \$1.95
Table 4: Effect of the number of selected keywords on coarse reranking: performance, token usage and cost.
Our experiments reveal a trade-off between performance and cost when varying the number of keywords. In general, increasing the number of selected keywords leads to a linear increase in token cost, while also improving coarse reranking performance. However, we observe diminishing returns beyond 10 keywords. This is expected, as keywords are included in order of embedding similarity to the query—those added later tend to be less relevant and contribute weaker relevance signals.
Fine-Grained Pool Size.
In the main experiments, for fine-grained reranking, we select the top 20 candidates from the coarse reranking stage to align with other reranking baselines. Here, we explore the impact of varying this candidate selection size. Specifically, we evaluate performance when selecting 5, 10, 20, 40, 60, 80, and 100 candidates for fine-grained reranking. The results are shown in Table 5.
Pool Size N@10 M@10 R@10 Token Usage Cost
5 docs 43.0 41.3 46.7 0.25M \$0.10
10 docs 44.3 42.8 46.7 0.50M \$0.20
20 docs 46.3 44.7 49.4 1.01M \$0.40
40 docs 45.4 43.9 48.0 2.01M \$0.81
60 docs 45.0 43.4 48.3 3.02M \$1.21
80 docs 45.3 43.8 48.0 4.03M \$1.61
100 docs 46.0 44.4 49.1 5.03M \$2.01
Table 5: Effect of the fine-grained candidate pool size: performance, token usage and cost.
Experimental results show that increasing the fine-grained reranking candidate pool from 0 to 20 documents yields a favorable trade-off between token cost and performance. However, performance saturates beyond 20 candidates and remains stable up to 100 documents, despite the growth of token consumption.
In theory, if an LLM’s long-context capability were fully effective, performance should continue to improve as more candidates are included. In practice, however, the observed plateau suggests that the effective context length is often much shorter than the theoretical context limit, which is also mentioned in previous studies (Hsieh et al., 2024; Kuratov et al., 2024).
5 Related Work
Classical Document Reranking.
Document Reranking (Karpukhin et al., 2020; Ren et al., 2021, 2023) is a critical component in information retrieval (IR) (Baeza-Yates et al., 1999; Singhal et al., 2001), especially when high retrieval precision is required and the first-stage retriever alone is insufficient. Early reranking methods were typically lexical and probabilistic (Salton et al., 1975; Robertson and Jones, 1976), relying on term overlap between the query and document to adjust relevance scores. However, these approaches were limited by reliance on exact matching and failed to capture in-depth semantics (Karpukhin et al., 2020; Gao et al., 2021a).
With advances in deep learning, neural reranking methods (Guo et al., 2019; Trabelsi et al., 2021), particularly those based on learning-to-rank (Cao et al., 2007; Liu, 2010) frameworks became prevalent. These models overcome previous limitations, capturing semantic similarity beyond surface-level term matches.
The emergence of pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Reimers and Gurevych, 2019) has further transformed the field of document reranking (Nogueira et al., 2019a; Yates et al., 2021). Cross-encoder-based rerankers (Litschko et al., 2022), in particular, have demonstrated strong performance by modeling query-document relevance scores and optimizing pointwise, pairwise, or listwise loss functions (Zhuang et al., 2023).
LLM-Based Listwise Document Reranking.
Recently, large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) have revolutionized the field of NLP. With strong general knowledge and instruction-following (Ouyang et al., 2022) capabilities, LLMs offer a new paradigm for listwise document reranking. A classic proposal (Sun et al., 2023a) involves placing the query and a list of candidate documents into the context window and prompting the LLM to generate a reordered ranking. This zero-shot strategy has been shown to outperform traditional reranking baselines (Sun et al., 2023a; Ma et al., 2023).
After that, LLM listwise reranking has been extended in multiple directions. Works like RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b) represent early explorations of open-source reranking LLMs. For inference time efficiency, FIRST (Gangi Reddy et al., 2024) modify listwise reranking by using only the first token decoding. RankMistral (Liu et al., 2024c) adopts long-context training to help LLMs adapt to larger document lists in a single input. Test-time scaling (Xia et al., 2025) also emerged as a common enhancement. For example, the sliding window strategy (Sun et al., 2023a; Ma et al., 2023) partitions the candidate list into overlapping chunks, enabling broader coverage within limited context. Building on this, Tang et al. (2024) enhance reranking quality through permutation consistency. ScaLR (Ren et al., 2024) further designs self-calibration to improve consistency across windows.
Notably, two prior studies (Liu et al., 2024b; Li et al., 2024) also identified limitations of full-document representations, but their analyses are on general domain. Moreover, the alternative representations they propose are fundamentally different from ours—one (Liu et al., 2024b) adopts non-natural language embeddings, and the other (Li et al., 2024) relies on document chunking.
To the best of our knowledge, our work is the first to highlight the unique limitations of full-document representations in scientific retrieval. We are also the first to leverage IE-based features as document representations for reranking. This makes our contribution novel and significant within the reranking literature.
6 Conclusion
In this work, we first identify fundamental challenges of standard listwise reranking in scientific retrieval. On the one hand, the first-stage retriever often fails to rank truly relevant documents high due to the lack of domain-specific training data. On the other hand, existing LLM-based listwise reranking methods typically operate over full-document inputs, which are substantial in terms of token usage and therefore limited to a small number of candidates. As a result, when the first-stage retrieval is suboptimal for scientific retrieval, the performance of standard reranking method is massively restricted.
To address this, we propose CoRank, a training-free and model-agnostic reranking framework for scientific retrieval. CoRank employs compact semantic features for coarse reranking. This allows the LLM to consider a broader range of candidates within its context window. A subsequent fine-grained reranking stage then refines the top results using full-text inputs. Experiments on two scientific benchmarks show that CoRank significantly improves reranking performance.
Overall, we are the first to explore semantic features as compact document representations for LLM reranking in scientific retrieval. Our results highlight the value of semantic feature extraction as a pre-analysis step in LLM listwise reranking in the scientific domain.
Limitations
Feature Extraction Quality.
In our offline preprocessing stage, we apply zero-shot information extraction using LLMs to obtain semantic features such as categories for each document. This is a relatively simple approach. While it has proven effective based on both quantitative results and case studies, there is still potential room for improvement. For instance, one could explore using more specialized models or introducing multi-turn feedback mechanisms to enhance the quality of extracted features. That said, since the primary focus of our work is to address the reranking bottleneck in scientific retrieval, we did not dedicate significant effort to modifiying the IE component, especially observing that the basic extraction already yielded strong performance.
Applicability Across Domains.
This work focuses exclusively on scientific retrieval. The motivation and assumptions behind our method are specifically tailored to the characteristics of this domain, such as long-tail terminology and the limitations of first-stage retrievers. We do not evaluate our method in general-domain settings. The potential effectiveness and limitations of our method outside the scientific domain remain untested and are left for future work.
Ethics Statement
We conduct our experiments on widely recognized and publicly available scientific retrieval datasets. Our proposed method and findings do not pose any foreseeable harm to individuals or groups. Overall, we do not anticipate any significant ethical concerns with this work.
Appendix A Prompt Templates
In the preprocessing stage, our method uses Qwen3-8B-Instruct (Qwen et al., 2024) to perform zero-shot information extraction (Wei et al., 2023; Sainz et al., 2024), instructing the model to extract document-level semantic features such as categories (Sun et al., 2023b; Zhang et al., 2024), sections (Zhou et al., 2023), and keywords (Rose et al., 2010; Lee et al., 2023). During inference, we provide a zero-shot, model-agnostic reranking framework, and evaluate its performance on Qwen3-32B-Instruct (Qwen et al., 2024), Gemini 2.0 Flash (Google DeepMind, 2025), and GPT-4.1-mini (OpenAI, 2024).
Here, we include the prompts used for different types of zero-shot document-level information extraction and model-agnostic reranking.
A.1 Extracting Categories
Prompt Template
Please analyze this document for its topic and categories:
{document} Provide a comprehensive analysis that includes: 1. The broad category (coarse-grained) this document belongs to 2. The specific category (fine-grained) within that broad category 3. A concise, title-like description of the document’s topic Deliver the analysis in one concise paragraph.
A.2 Extracting Sections
Prompt Template
Identify 3-8 logical sections that would effectively organize this document’s content:
{document} Generate appropriate subtitle-style headings for each section that would help structure the document. Sections should be comprehensive and cover the full scope of the content.
A.3 Extracting Keywords
Prompt Template
Extract a comprehensive list of at least 30 diverse keywords and concepts from this document:
{document} Generate as many diverse, relevant keywords and concepts as possible. Include both specific terms and broader conceptual themes.
A.4 Generating Pseudo Queries
Prompt Template
Generate 20 diverse search queries that users might enter to find this document:
{document} Create different types of queries that cover various aspects of the document content. Queries should be diverse in wording, length, and specificity.
A.5 Reranking
Prompt Template
You are an LLM reranker, an intelligent assistant that can rank passages based on their relevancy to the query.
I will provide you with {num} passages (either represented by full text, previous user query, keywords or structured analysis), each indicated by a numerical identifier []. Rank the passages based on their relevance to the search query: {query}. [1]{passage} [n]{passage} Search Query: {query}. Rank the {num} passages above based on their relevance to the search query. All the passages should be in descending order of relevance. The output format should be [passage_id] > [passage_id] > …, (If the full list is very long, generate at least 10) e.g., [4] > [2] > … Only respond with the ranking results, do not say any word or explain.
Appendix B Case Studies
B.1 Distribution of Positives and Negatives
While metrics like nDCG@10 quantify the overall reranking quality, they do not indicate which specific reranking judgement contribute to the improvements. To better understand where our performance gains and limitations come from, we perform a qualitative analysis of how the position of ground-truth documents changes in the ranked list with or without CoRank.
We compare CoRank with the vanilla zero-shot listwise reranker on LitSearch (Ajith et al., 2024) with GPT-4.1-mini (OpenAI, 2024). Based on the rank position of the ground-truth documents and whether they fall within the top-10 (i.e., Recall@10), we obtain the distribution of positive cases and negative cases.
Figure 4: Distribution of Recall@10 positive and negative in Vanilla listwise reranking and CoRank.
Analysis on Shared Positives.
As shown in figure 4, CoRank successfully covers the majority of positive cases (69 out of 71) of the vanilla reranker. This is expected, as the fine-grained reranking stage in CoRank inherits the full-text completeness in document representation of the vanilla listwise reranking.
Analysis on Exclusive Positives.
At the same time, CoRank exclusively attains 58 positive cases. Analyzing these cases, we find that most (57 out of 58) were missed by the vanilla reranker because their first-stage retrieval ranks were too low (i.e., outside the top 20). This supports our analysis in Section 2.2 on the key challenge in scientific-domain reranking: truly relevant documents may be excluded from reranking due to suboptimal first-stage retrieval results. CoRank mitigates this issue with the compactness of feature-based representations in the coarse reranking stage, allowing more candidates to be considered and increasing the likelihood of recovering relevant documents.
Next, we present full instances of both positive and negative examples, along with their ranks in the first-stage retrieval, coarse reranking stage, and fine-grained reranking stage. This allows us to analyze the underlying causes of success or failure in each case.
B.2 Positive Example
This example clearly demonstrates the advantages of our method. The ground-truth document in this case involves long-tailed scientific concepts such as CGA and VAE, which are not well captured by the first-stage retriever. As a result, it is ranked only 92nd in the initial retrieval, making it inaccessible to vanilla rerankers that operate on the top 20 candidates. Even with the sliding window strategy (Sun et al., 2023a; Ma et al., 2023), the document would need to consistently win across 9 windows to enter the top-10, highlighting the limitations of existing reranking methods in scientific domains, as discussed earlier. In contrast, our method successfully ranks this document as the top candidate during the coarse reranking stage. This illustrates two key points: (1) our feature-based representation is highly compact, allowing us to expand the reranking pool to 200 candidates and include low-ranked but relevant documents such as this one; and (2) it is also sufficiently informative, enabling the reranker to recognize and promote the document directly to the top of the list once included.
Query
Can you recommend a foundational paper that provides a scalable framework for generating English sentences with controllable semantic and syntactic attributes for the purpose of augmenting datasets in NLP tasks?
Ground Truth Document Control, Generate, Augment: A Scalable Framework for Multi-Attribute Text Generation: We introduce CGA, a conditional VAE architecture, to control, generate, and augment text. CGA is able to generate natural English sentences controlling multiple semantic and syntactic attributes by combining adversarial learning with a context-aware loss and a cyclical word dropout routine. We demonstrate the value of the individual model components in an ablation study. The scalability of our approach is ensured through a single discriminator, independently of the number of attributes. We show high quality, diversity and attribute control in the generated sentences through a series of automatic and human assessments. As the main application of our work, we test the potential of this new NLG model in a data augmentation scenario. In a downstream NLP task, the sentences generated by our CGA model show significant improvements over a strong baseline, and a classification performance often comparable to adding same amount of additional real data.
First-Stage Retrieval Ranking 92 of 14256
Compact Representation Natural Language Processing (NLP) -> Text Generation and Neural Machine Translation -> Conditional VAE-Based Framework for Controllable and Scalable Multi-Attribute Text Generation with Applications in Data Augmentation: CGA for Data Augmentation in NLP Tasks (Text Generation, Multi-Attribute Control, Data Augmentation, Semantic Attributes)
Coarse Reranking 1 of 200
Fine-grained Reranking 1 of 20
B.3 Negative Example
The example above illustrates a case where our method fails to rank the ground-truth document highly—not due to representation quality, but because the document was ranked extremely low (319th) by the first-stage retriever. At such a low position, it is not visible to vanilla reranking methods, nor to those with extended context windows such as the sliding window strategy (Sun et al., 2023a; Ma et al., 2023) or our Method.
As a result, even though our compact representation accurately captures the document’s semantic content, it never enters the reranking process and thus has no opportunity to be promoted to a higher position.
Query
Are there any studies that explore post-hoc techniques for hallucination detection at both the token- and sentence-level in neural sequence generation tasks?
Ground Truth Document Detecting Hallucinated Content in Conditional Neural Sequence Generation: Neural sequence models can generate highly fluent sentences, but recent studies have also shown that they are also prone to hallucinate additional content not supported by the input. These variety of fluent but wrong outputs are particularly problematic, as it will not be possible for users to tell they are being presented incorrect content. To detect these errors, we propose a task to predict whether each token in the output sequence is hallucinated (not contained in the input) and collect new manually annotated evaluation sets for this task. We also introduce a method for learning to detect hallucinations using pretrained language models fine tuned on synthetic data that includes automatically inserted hallucinations Experiments on machine translation (MT) and abstractive summarization demonstrate that our proposed approach consistently outperforms strong baselines on all benchmark datasets. We further demonstrate how to use the token-level hallucination labels to define a fine-grained loss over the target sequence in low-resource MT and achieve significant improvements over strong baseline methods.We also apply our method to word-level quality estimation for MT and show its effectiveness in both supervised and unsupervised settings 1.
Compact Representation Natural Language Processing -> Hallucination Detection in Neural Sequence Generation -> Token-Level Hallucination Detection in Conditional Text Generation: detecting hallucination using token-level classification and pretrained language models (Hallucination Detection, Neural Sequence Generation, Token-Level Classification, Machine Translation (MT))
First-Stage Retrieval Ranking 319 of 14256
Coarse Reranking N/A
Fine-grained Reranking N/A
Appendix C Proprietary Model API Cost
As part of our evaluation of different reranking methods, we tested two commercial LLMs: Gemini 2.0 Flash (Google DeepMind, 2025) and GPT-4.1-mini (OpenAI, 2024). Here, we report the API costs incurred in the main experiments when querying these models. The costs associated with experiments in the Further Analysis section have already been presented as part of the token efficiency comparison.
Model Price Setting # Token Cost (\$)
[…]
Total Cost 29.61M \$2.96
[…]
Total Cost 29.61M \$11.84
Table 6: Reranking token usage and cost across different settings for Gemini 2.0 Flash and GPT-4.1-mini (OpenAI, 2024). Prices are per 1 million tokens.
As shown in Table 6, across all different settings in the main experiments, we spent \$2.96 on Gemini-2.0-Flash (Google DeepMind, 2025) and \$11.84 on GPT-4.1-mini (OpenAI, 2024), resulting in a total API cost of \$14.80.
Appendix D Future Work
In the previous section, we identified two main limitations of our current approach. These suggest promising directions for future research that could further improve the generality and effectiveness of our method.
D.1 Improving Feature Extraction
Our current pipeline uses zero-shot information extraction (Wei et al., 2023) with general-purpose LLMs (Qwen et al., 2024) to generate semantic features such as categories (Sun et al., 2023b; Zhang et al., 2024) and keywords (Rose et al., 2010; Lee et al., 2023). While this approach is simple and has proven effective, future work could explore more advanced techniques to enhance extraction quality. For example, using domain-specific LLMs (Taylor et al., 2022; Xie et al., 2023) or introducing multi-turn feedback (Shinn et al., 2023) may yield richer and more accurate representations, further improving downstream reranking performance.
D.2 Extending to General Domains
Our method is designed specifically for scientific retrieval (Lawrence et al., 1999; White et al., 2009), where challenges like long-tail terminology (Kang et al., 2024a) and limited first-stage recall (Kim et al., 2023) are particularly severe. An important direction for future work is to investigate the applicability of our framework in general-domain retrieval settings (Zhu et al., 2023; Ai et al., 2023). This includes evaluating how well the feature-based representation and the hybrid reranking strategy transfer beyond the scientific context, and identifying any necessary adaptations.

### v2

@@
-LLM-Based Compact Reranking with Document Features for Scientific Retrieval
-Abstract
-Scientific retrieval is essential for advancing academic discovery.
-Within this process, document reranking plays a critical role by refining first-stage retrieval results.
-However, standard LLM listwise reranking faces unique challenges in the scientific domain.
+CoRank: LLM-Based Compact Reranking with Document Features for Scientific Retrieval
+Abstract.
+Scientific retrieval is essential for advancing scientific knowledge discovery.
+Within this process, document reranking plays a critical role in refining first-stage retrieval results.
+However, standard LLM listwise reranking faces challenges in the scientific domain.
 First-stage retrieval is often suboptimal in the scientific domain, so relevant documents are ranked lower.
-Moreover, conventional listwise reranking inputs the full text of candidates into the context window, limiting the number of candidates that can be considered.
+Meanwhile, conventional listwise reranking places the full text of candidates into the context window, limiting the number of candidates that can be considered.
 As a result, many relevant documents are excluded before reranking, constraining overall retrieval performance.
-To address these challenges, we explore compact document representations based on semantic features (e.g., categories, sections and keywords) and propose CoRank, a training-free, model-agnostic reranking framework for scientific retrieval.
-It presents a three-stage solution: (i) offline extraction of document-level features, (ii) coarse reranking using these compact representations, and (iii) fine-grained reranking on full texts of the top candidates from (ii).
-This hybrid design provides a high-level abstraction of document semantics, expands candidate coverage, and retains critical details required for precise ranking.
-Experiments on LitSearch and CSFCube show that CoRank significantly improves reranking performance across different LLM backbones (nDCG@10 from $32.0$ to $39.7$).
-Overall, these results highlight the value of information extraction for reranking in scientific retrieval.
-Code will be publicly available.
-LLM-Based Compact Reranking with Document Features for Scientific Retrieval
-Figure 1: Instead of directly reranking full documents, we extract features, rerank a larger candidate pool with them, and finally refine the shortlist with full documents.
-1 Introduction
-Scientific retrieval (Lawrence et al., 1999; White et al., 2009) is crucial for scientific discovery.
-While current retrievers are effective at retrieving broadly relevant scientific papers, they often struggle to differentiate between documents covering similar topics (Sciavolino et al., 2021; Liu et al., 2021), making fine-grained relevance estimation essential.
+To address these challenges, we explore semantic-feature-based compact document representations (e.g., categories, sections, and keywords) and propose CoRank, a training-free, model-agnostic reranking framework for scientific retrieval.
+It presents a three-stage solution: (i) offline extraction of document features, (ii) coarse-grained reranking using these compact representations, and (iii) fine-grained reranking on full texts of the top candidates from (ii).
+This integrated process addresses suboptimal first-stage retrieval: Compact representations allow more documents to fit within the context window, improving candidate set coverage, while the final fine-grained ranking ensures a more accurate ordering.
+Experiments on 5 academic retrieval datasets show that CoRank significantly improves reranking performance across different LLM backbones (average nDCG@10 from $50.6$ to $55.5$).
+Overall, these results underscore the synergistic interaction between information extraction and information retrieval, demonstrating how structured semantic features can enhance reranking in the scientific domain.
+Information Retrieval; Scientific Document Search; Large Language Models; LLM-Based Reranking
+conference: ACM International Conference on Web Search and Data Mining; Feb 22–26, 2026; Boise, USAjournalyear: 2026ccs: Information systems Retrieval models and rankingccs: Computing methodologies Natural language processing
+1. Introduction
+Figure 1. We extract semantic features, rerank a larger candidate pool with them, and finally refine the shortlist with full text instead of directly reranking with full text.
+[…]
+Flow diagram with three stages: (i) offline extraction of document-level features; (ii) coarse-grained reranking over a large candidate set using compact feature vectors; (iii) fine-grained reranking of the shortlisted candidates using full-text representation.
+Arrows indicate the processing order from (i) to (iii).
+Scientific retrieval (Lawrence et al., 1999; White et al., 2009) is crucial for scientific knowledge discovery.
+While current retrievers are effective at retrieving broadly relevant scientific papers, they often struggle to differentiate among documents covering similar topics (Sciavolino et al., 2021; Liu et al., 2021), failing to make fine-grained relevance estimation.
 Therefore, the reranking stage (Carbonell and Goldstein-Stewart, 1998; Kurland and Lee, 2005) is particularly important (Gao et al., 2021b), as it refines the first-stage retrieval to better distinguish between closely ranked documents.
-Recently, large language models (LLMs) (Grattafiori et al., 2024; OpenAI et al., 2024) have significantly advanced document reranking, particularly through their application to listwise reranking (Sun et al., 2023a; Ma et al., 2023).
+Recently, Large Language Models (LLMs) (Grattafiori et al., 2024; OpenAI et al., 2024) have significantly advanced the field of document reranking, particularly through their application in listwise reranking (Sun et al., 2023b; Ma et al., 2023).
 In this setting, LLMs jointly assess a set of retrieved candidates within their context window and generate a reordered ranking based on their relevance.
-Previous works (Pradeep et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c; Ren et al., 2024) show that LLM-based listwise rerankers outperform prior embedding-based approaches (Nogueira et al., 2019b, 2020) on benchmarks like BEIR (Thakur et al., 2021).
-The standard practice in LLM listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) is to input the full text of each candidate document into the context window.
-However, this approach faces key limitations in scientific retrieval.
-In the scientific domain, retrievers often show limited performance (Kim et al., 2023; Kang et al., 2024a, b), so truly relevant documents may not rank high enough in the first-stage retrieval.
-Moreover, due to the substantial token overhead of full text representation and finite context length, rerankers can only operate on a limited number (typically 20 per prompt) (Sun et al., 2023a; Ma et al., 2023) of candidates.
-As a result, when first-stage retrieval is suboptimal, the reranking performance is inherently constrained (Reddy et al., 2023).
-To address these limitations, we explore an alternative document representation that is both compact and informative.
-Instead of using full text, we investigate an information extraction (IE)–based approach (Niklaus et al., 2018; Zhou et al., 2022; Liu et al., 2022), representing each scientific paper using high-level features such as categories, sections, and keywords.
-We find that this representation significantly improves per-document token efficiency, enabling a larger pool of candidates to fit within the LLM’s context window.
-This, in turn, makes the reranking process more robust to suboptimal first-stage scientific retrieval results.
-It also simplifies the input by filtering out irrelevant details, making it easier for LLMs to interpret.
-However, since IE cannot guarantee complete coverage of all relevant content, full-text inputs still serve as a valuable complement when fine-grained relevance comparisons are required.
-Motivated by these insights, we propose CoRank, a training-free, model-agnostic reranking framework for science retrieval.
-CoRank consists of three stages: (i) Offline Preprocessing: extract high-level semantic features like categories and keywords from unstructured scientific documents; (ii) Coarse Reranking: use these compact representations for an initial ranking and select a subset of top candidates; (iii) Fine-grained Reranking: rerank the top candidates using full scientific documents to recover the potentially missing details during information extraction.
-Our design enhances the robustness against suboptimal first-stage retrieval with compact feature representation and maximizes the overall effectiveness with final full text reranking.
-We evaluate CoRank on two scientific retrieval benchmarks: LitSearch (Ajith et al., 2024) and CSFCube (Mysore et al., 2021).
-Empirical experiments show that across various LLM backbones, CoRank achieves an absolute improvement of $+11.5$ nDCG@10 without the sliding window strategy, and retains a $+3.9$ gain with it.
-These results demonstrate significant effectiveness of our method for reranking in scientific retrieval.
-Overall, our contributions are threefold:
+Previous works (Pradeep et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024b; Ren et al., 2024) show that LLM-based listwise rerankers outperform prior embedding-based approaches (Nogueira et al., 2019, 2020) on multiple evaluation datasets (Thakur et al., 2021).
+The standard practice in LLM listwise reranking (Sun et al., 2023b; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024b) is to input the full text of each candidate document into the context window.
+In scientific retrieval, however, this approach faces key limitations.
+In the scientific domain, first-stage retrievers often show limited performance (Kim et al., 2023; Kang et al., 2024a, b), so truly relevant documents may not rank high enough in the first-stage retrieval.
+Moreover, due to the substantial token overhead of full text representation and finite context length, rerankers can only operate on a limited number (typically 20 per prompt) (Sun et al., 2023b; Ma et al., 2023) of candidates.
+Consequently, when the first-stage retrieval is suboptimal, many relevant documents are excluded from reranking consideration, which inherently constrains overall performance (Reddy et al., 2023).
+A straightforward solution is to expand the context window of LLM rerankers to handle more candidates.
+This can be done using long-context LMs (Huang et al., 2023; Pawar et al., 2024) or test-time scaling strategies (Snell et al., 2024; Xia et al., 2025), such as multi-round reranking (Sun et al., 2023b; Ma et al., 2023).
+However, these methods have notable drawbacks: long-context LMs require additional training (Liu et al., 2024b), suffer from “loss in the middle” problem (Liu et al., 2024a; Hsieh et al., 2024), and incur high computational costs (Liu et al., 2025b); sliding windows similarly increase inference latency and overhead (Sun et al., 2023b; Ma et al., 2023).
+To better address the limitations, we explore compressing each document’s representation by abstracting high-level features (e.g., categories, keywords), enabling more documents to fit within a fixed context window.
+This achieves a similar effect to window expansion but with greater efficiency.
+We propose CoRank, a zero-shot, model-agnostic reranking framework for science retrieval as exhibited in Figure 1.
+CoRank consists of three stages: (i) Offline Preprocessing: extract high-level semantic features like categories and keywords from unstructured scientific documents; (ii) Coarse-grained Reranking: use these compact representations for an initial ranking and select a subset of top candidates; (iii) Fine-grained Reranking: rerank the top candidates using full scientific documents to recover the potentially missing details during information extraction.
+This integrated reranking framework aims to address the suboptimal first-stage retrieval.
+Compact representations allow more documents to fit within the context window, improving candidate set coverage, while the final fine-grained ranking ensures a more accurate ordering.
+We comprehensively evaluate CoRank on five diverse scientific retrieval datasets spanning computer science (Ajith et al., 2024; Mysore et al., 2021), biomedical (Boteva et al., 2016), and clinical (Wadden et al., 2020; Voorhees et al., 2021) domains.
+Our experiments cover a diversity of LLM backbones, including both proprietary (OpenAI et al., 2024; Google DeepMind, 2025) and open-source (Qwen et al., 2024; Jiang et al., 2023; Grattafiori et al., 2024) models.
+Results show that CoRank consistently outperforms existing LLM-based zero-shot reranking baselines, achieving an average absolute improvement of +4.9+4.9 nDCG@10.
+Notably, our method delivers strong gains even in challenging settings with suboptimal first-stage retrieval, while also maintaining high efficiency by reducing token consumption compared to the widely adopted sliding window strategy (Sun et al., 2023b).
+Extensive ablation studies and further analyses validate the contribution of each component and demonstrate the robustness of our method.
   1. # 1
-    We are the first to reveal the unique limitations of LLM listwise reranking in scientific retrieval.
+    We reveal the limitations of LLM listwise reranking in scientific retrieval.
   2. # 2
-    We are the first to explore semantic features as compact representations in reranking and propose CoRank, a framework based on this design.
+    We explore semantic features for document representation in reranking and propose CoRank, a novel zero-shot reranking framework.
   3. # 3
-    We show that CoRank significantly improves reranking performance in scientific retrieval.
-These findings underscore the value of compact feature extraction as a pre-analysis step for improving the reranking performance in scientific retrieval.
-Figure 2: Overview of our feature extraction pipeline: from unstructured documents, we apply zero-shot LLM information extraction to obtain document features including categories, sections, pseudo queries, and keywords, which are then combined into compact representations.
-2 Preliminary Analysis
-In this section, we first define listwise reranking, then show the limitations of current full text-based methods and explore compact and informative document representation options for reranking in the scientific domain.
-2.1 Problem Definition
+    We show that CoRank significantly improves reranking performance of LLM-based listwise approach in scientific retrieval.
+These findings underscore how semantic information extraction and retrieval interact, with pre-analysis of document features enabling more effective reranking in scientific search.
+2. Related Work
+2.1. Classical Document Reranking
+Document reranking (Karpukhin et al., 2020; Ren et al., 2021, 2023) is a critical component in information retrieval (IR)  (Baeza-Yates et al., 1999; Singhal et al., 2001), especially when high retrieval precision is required and the first-stage retriever alone is insufficient.
+Early reranking methods were lexical and probabilistic  (Salton et al., 1975; Robertson and Jones, 1976).
+Like early sparse retrieval methods (Baeza-Yates et al., 1999; Robertson et al., 2009), they relied on term overlap between the query and document to adjust relevance scores.
+However, unlike initial retrieval, rerankers revisit a smaller set of candidate documents returned by the first stage and apply more sophisticated weighting or adjustment techniques to refine the ranking.
+Nevertheless, these approaches were limited by their reliance on exact matching and failed to capture in-depth semantics (Karpukhin et al., 2020; Gao et al., 2021a).
+With advances in deep learning, neural reranking methods (Guo et al., 2019; Trabelsi et al., 2021), particularly those based on learning-to-rank (Cao et al., 2007; Liu et al., 2009) frameworks, became prevalent.
+These models overcome previous limitations, capturing semantic similarity beyond surface-level term matches.
+The emergence of pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Reimers and Gurevych, 2019) further transformed document reranking (Nogueira and Lin, 2019; Yates et al., 2021).
+Cross-encoder-based rerankers (Litschko et al., 2022), in particular, have demonstrated strong performance by modeling query-document relevance scores and optimizing pointwise, pairwise, or listwise loss (Zhuang et al., 2023).
+Unlike bi-encoder architectures used in dense retrieval (Karpukhin et al., 2020), cross-encoders jointly encode both the query and each candidate document, allowing the model to directly capture rich dependencies between queries and documents from first-stage retrieval.
+Figure 2.
+Overview of our feature extraction pipeline: from unstructured documents, we apply zero-shot LLM information extraction to obtain document features including categories, sections, pseudo queries, and keywords, which are then combined into compact representations.
+[…]
+Flow diagram of the document representation pipeline.
+Left: an unstructured scientific document enters the system.
+Middle: a zero-shot LLM information extraction step produces four feature sets—categories (three-level path), section titles, pseudo queries, and keywords.
+An adaptive selection module keeps the most relevant items (e.g., one section, one pseudo query, and top-k keywords).
+Right: the selected items are concatenated into a compact representation that is passed to the reranker.
+Arrows indicate the order: document → extraction → selection → compact representation → reranking.
+2.2. LLM-Based Listwise Document Reranking
+Recently, large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) have revolutionized the field of NLP.
+With strong general knowledge and instruction-following (Ouyang et al., 2022) capabilities, LLMs offer a new paradigm for listwise document reranking.
+A classic proposal (Sun et al., 2023b) involves placing the query and a list of candidate documents into the context window and prompting the LLM to generate a reordered ranking.
+This zero-shot strategy has been shown to outperform the traditional reranking baselines (Sun et al., 2023b; Ma et al., 2023).
+After the introduction of LLM-based listwise reranking, research has rapidly extended in several directions.
+For model development, works such as RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b) pioneered open-source reranking LLMs, while RankMistral (Liu et al., 2024b) leveraged long-context training to adapt LLMs for larger document lists in a single input.
+For context-length adaptation, Sliding window strategies (Sun et al., 2023b; Ma et al., 2023) partition candidate lists into overlapping chunks for broader coverage, addressing the limited context windows of LLMs.
+For inference efficiency, FIRST (Gangi Reddy et al., 2024) improves the speed of listwise reranking by restricting the process to first-token decoding.
+Liu et al.
+(2025a) and Li et al.
+(2024a) use passage embeddings and chunked text blocks to represent documents reducing the token cost per document.
+From the perspective of self-consistency, ScaLR (Ren et al., 2024) introduces self-calibration to improve consistency across windows.
+Meanwhile, Tang et al.
+(2024) enhance reranking quality through permutation consistency.
+Since our method operates on document representations in a plug-and-play fashion, it is orthogonal to most previous baselines and can be applied jointly with them (we illustrate this with the sliding window strategy in Section 5).
+Notably, two prior studies (Liu et al., 2025a; Li et al., 2024a) have also identified limitations of full-document representations, but their approaches are fundamentally different from ours: one (Liu et al., 2025a) adopts non-natural language embeddings, while the other (Li et al., 2024a) merely relies on document chunking.
+Furthermore, their analyses focus on the general domain, whereas ours is conducted in the scientific domain.
+Overall, our work highlights the limitations of full-document representations for listwise reranking in scientific retrieval.
+And we leverage IE-based features as document representations for reranking.
+This makes our contribution novel and significant within the reranking literature.
+Figure 3.
+Token efficiency and performance comparison across different document representations.
+(a) Per-document token length distribution.
+(b) nDCG@10 scores for different numbers of documents in the context window of single LLM input (c) Context token overhead to reach equal reranking performance.
+[…]
+Composite figure with three plots.
+(a) Violin/box plots show token counts per document: full text averages around 200 tokens, with a long tail past 5,000; four feature-based forms (Form 1–4) are far shorter, roughly 10–50 tokens each.
+(b) Line chart of nDCG@10 as candidate-set size rises from 10 to 200.
+Full text reranking leads up to  80 documents, but exceeds the 32k-token limit thereafter; Form 4 continues improving and overtakes full text beyond 100 candidates.
+(c) Bar chart of total context tokens required to reach 70%, 85%, 100%, 115%, and 130% of the 20-doc full text baseline.
+Full text needs ≈3,500 tokens for 100% performance, whereas Form 4 reaches the same level with ≈1,000 tokens and achieves 130% performance with fewer tokens than full text at 100%.
+The visual message: feature-based representations are dramatically more token-efficient, allowing more documents per prompt and equal or better reranking quality.
+3. Preliminary Analysis
+In this section, we first define the task of listwise reranking, then show the limitations of current full-document-based methods and explore compact and informative document representation options for reranking in the scientific domain.
+3.1. Problem Definition
 Reranking Input.
@@
 This candidate list $\mathcal{C}$ is then passed to a reranking model, which aims to reorder the documents such that more relevant ones are ranked higher.
-Reranking Objective.
-The goal of document reranking is to find a permutation $\sigma:\{1,\dots,m\}\!\to\!\{1,\dots,m\}$ whose application
+Reranking Objective. The goal of document reranking is to find a permutation $\sigma:\{1,\dots,m\}\!\to\!\{1,\dots,m\}$ whose application
 $\mathcal{C}^{\prime}\;=\;[\,p_{\sigma(1)},\,p_{\sigma(2)},\dots,p_{\sigma(m)}\,]$
 orders documents in more accurate descending relevance to $q$.
-The quality is often (Sun et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c) evaluated using top-$k$ metrics such as nDCG@10, with small $k$ values that stress accuracy at the top of the ranked list.
+The quality is often (Sun et al., 2023b; Gangi Reddy et al., 2024; Liu et al., 2024b) evaluated using top-$k$ metrics such as nDCG, with small $k$ values that stress accuracy at the top of the ranked list.
 LLM Listwise Reranking.
-Large language models can inspect $m$ candidate documents in their context window and generate a permutation that reflects their relevance ordering. Formally, the model acts as follows:
-$\sigma_{\text{LLM}}\;=\;\mathrm{LLM}(q,\,\mathcal{C})\;\in\;S_{m},$
-where $S_{m}$ is the set of all permutations on $\{1,\dots,m\}$. Applying this permutation to the candidate list yields
+Large language models can inspect $m$ candidate documents in their context window and generate a permutation that reflects their relevance ordering.
+Formally, the model acts as follows:
+$\sigma_{\text{LLM}}\;=\;\mathrm{LLM}(q,\,\mathcal{C})$
+Applying this permutation to the candidate list yields
 $\mathcal{C}^{\prime}\;=\;\bigl{[}p_{\sigma_{\text{LLM}}(1)},\;p_{\sigma_{\text{LLM}}(2)},\;\dots,\;p_{\sigma_{\text{LLM}}(m)}\bigr{]}.$
-2.2 Current Limitations
-The standard approach in LLM-based listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) feeds the full text of each candidate into the model’s context window.
-While this allows LLMs to capture the complete content of each document, it incurs some limitations, especially in scientific retrieval.
+3.2. Current Limitations
+The standard approach in LLM-based listwise reranking (Sun et al., 2023b; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024b) feeds the full text of each candidate into the model’s context window.
+While this allows LLMs to capture the complete content of each document, it faces some challenges in the scientific domain.
 Suboptimal First-Stage Retrieval.
-Unlike in general-domain scenarios, first-stage retrievers—whether sparse (Robertson et al., 2009; Nogueira et al., 2019a) or dense (Izacard et al., 2021; Wang et al., 2022)—struggle to generalize to the scientific domain (Thakur et al., 2021; Bonifacio et al., 2022).
+Unlike in general-domain scenarios, first-stage retrievers—whether sparse (Robertson et al., 2009; Nogueira and Lin, 2019) or dense (Izacard et al., 2021; Wang et al., 2022)—struggle to generalize to the scientific domain (Thakur et al., 2021; Bonifacio et al., 2022).
 This is due to the complexity of long-tail concepts (Kang et al., 2024b, 2025) and the lack of large-scale supervised training data (Bonifacio et al., 2022; Li et al., 2023).
@@
 Full text representation also introduces significant token overhead, with individual scientific papers often consuming hundreds or even thousands of tokens (Thakur et al., 2021; Ajith et al., 2024; Mysore et al., 2021).
-At the same time, LLMs have a limited effective context length (Dai et al., 2019; Xiong et al., 2024) and are known to suffer from issues such as positional bias for long inputs (Liu et al., 2024a; Tian et al., 2024).
-As a result, full-text rerankers are constrained to operate on a small number of candidates, typically 20 documents per input prompt (Sun et al., 2023a; Ma et al., 2023; Pradeep et al., 2023a, b).
+At the same time, LLMs have a limited effective context length (Dai et al., 2019; Xiong et al., 2024) and are known to suffer from issues such as positional bias for long inputs (Liu et al., 2024a; Tian et al., 2025).
+As a result, full text rerankers are constrained to operate on a small number of candidates, typically 20 documents per input prompt (Sun et al., 2023b; Ma et al., 2023; Pradeep et al., 2023a, b).
 When first-stage retrieval is suboptimal and reranking operates over a narrow candidate set, the overall reranking performance is significantly constrained.
-2.3 Semantic Features as Alternatives
-Given the limitations of full-text representations in scientific-domain reranking, we explore an alternative approach, which is based on document-level semantic features such as categories and keywords.
+3.3. Semantic Features as Alternatives
+3.3.1. Semantic Feature Construction
+Given the limitations of full text representations in scientific-domain reranking, we explore an alternative approach, which is based on document-level semantic features such as categories and keywords.
 The intuition is that these features are both significantly more concise and capable of preserving the core semantics.
@@
 Category.
-Categories (Sun et al., 2023b; Zhang et al., 2024) offer a high-level topical overview of scientific documents.
+Categories (Sun et al., 2023a; Zhang et al., 2025) offer a high-level topical overview of scientific documents.
 Each document is assigned a three-level hierarchical category path in the format {Category} → {Subcategory} → {Subsubcategory}, which provides a broad-to-specific classification on its topic.
@@
 This feature offers a unique query-aligned perspective of the document, simulating how the document might be retrieved in real-world scenarios.
-These features cover most common types of document-level IE and vary in granularity and style.
-All semantic features are extracted using LLM-based zero-shot information extraction.
-The prompt templates used for this extraction are provided in Appendix A.
+These four features represent the most common types of document-level information extraction (IE) (Niklaus et al., 2018; Wimalasuriya and Dou, 2010), varying in both granularity and style.
+All semantic features are extracted using LLM-based information extraction methods (Yuan et al., 2023; Zhang et al., 2023).
+We adopt zero-shot prompting with instruction-tuned models such as GPT-4.1 (OpenAI, 2024) and the Qwen model family (Qwen et al., 2024).
+This approach leverages the inherent knowledge of pretrained language models (Min et al., 2023; Li et al., 2024b) to extract information without requiring task-specific training (Yuan et al., 2023), while demonstrating strong generalization across domains and schema variations (Xu et al., 2025).
 Among these features, sections, keywords, and pseudo queries have multiple elements for each document.
-To reduce noise and focus on the most relevant content, we apply an adaptive selection strategy at inference time.
+To reduce noise and focus on the most relevant content, we apply an adaptive selection strategy during inference time.
 We compute dense embedding similarities between the query and each extracted element and retain only the most relevant: 5 keywords (from 30), 1 pseudo query (from 20), and 1 section (from 3).
 This improves content relevance while minimizing token overhead.
-Ablation of this strategy can be found in Section 4.2.
-Figure 3: Token efficiency and performance comparison across different document representations.
-(a) Per-document token lengths distribution.
-(b) nDCG@10 scores for different number of documents in the context window of single LLM input (c) Context token overhead to reach equal reranking performance.
 However, relying on a single feature often lacks representational power.
@@
     Category + Section + Keywords
-Ablation studies on the effect of each specific component can further be found in Section 4.2. The overview of representation construction is presented in Figure 2.
-2.4 Empirical Validation
+Ablation studies on the effect of each specific component can further be found in Section 5.2.2. The overview of representation construction is presented in Figure 2.
+3.3.2. Empirical Validation
 To evaluate the effectiveness of our feature-based representations, we conduct a series of experiments on the LitSearch (Ajith et al., 2024), using GPT-4.1-mini (OpenAI, 2024) as the reranking backbone.
-Semantic features are extracted using Qwen3-8B-Instruct (Qwen et al., 2024).
+Semantic features are extracted using Qwen3-8B (Qwen et al., 2024).
 Following Gangi Reddy et al.
@@
 We first evaluate the token cost of different representations.
-As shown in Figure 3 (a), full-text inputs consume around 200 tokens per document on average, with some exceeding 5,000 tokens, limiting the number of documents that can fit into the context window.
-In contrast, semantic feature-based representations (Forms 1-Form 4) are much more compact, each averaging between 10 and 50 tokens.
+As shown in Figure 3 (a), full text inputs consume around 200 tokens per document on average, with some exceeding 5,000 tokens, limiting the number of documents that can fit into the context window.
+In contrast, semantic-feature-based representations (Forms 1-4) are much more compact, each averaging between 10 and 50 tokens.
 Number of Documents in a Single Input.
@@
 As shown in Figure 3 (b), all representations improve as more documents are included.
-In the range of 0 to 80 candidates, full-text representations perform better than feature-based ones.
+In the range of 0 to 80 candidates, full text representations perform better than feature-based ones.
 This is expected, since full text retains complete document information, though at a much higher token cost.
-However, beyond 80 candidates, full-text inputs frequently exceed the context length limit.
+However, beyond 80 candidates, full text inputs frequently exceed the context length limit.
 In contrast, feature-based methods, especially Form 4, remain compact and can handle more candidates without issue, while still achieving comparable performance.
@@
 Finally, we examine how many context tokens are needed to achieve the same reranking performance across different representations.
-Using results from earlier experiments, we estimate the total token cost by combining the number of documents required with the average token count per document.
-We take full-text reranking with 20 documents as the 100% performance base, and compare how many tokens each method needs to reach 70%, 85%, 100%, 115%, and 130% of that level.
-As shown in Figure 3 (c), compact feature-based representations can match full-text performance with significantly fewer tokens.
+Using results from the previous two experiments, we estimate the total token cost by combining the number of documents required with the average token count per document.
+We take full text reranking with 20 documents as the 100% performance baseline, and compare how many tokens each method needs to reach 70%, 85%, 100%, 115%, and 130% of that level.
+As shown in Figure 3 (c), compact feature-based representations can match full text performance with significantly fewer tokens.
 For instance, reaching 100% performance requires around 3,500 tokens for full text, but only 1,000 or even a few hundred tokens for feature-based inputs.
-2.5 Discussion
-The results above demonstrate that feature-based document representations offer clear advantages over full-text inputs in reranking.
+3.3.3. Discussion
+The results above demonstrate that feature-based document representations offer clear advantages over full text inputs in reranking.
 They can express the core content of a document using far fewer tokens, allowing significantly more retrieved candidates to be included within the same token overhead.
 This is especially valuable for recovering relevant documents that were assigned lower scores by the first-stage retriever.
-Considering both token efficiency and reranking performance, we select Form 4 as our feature-based representation.
+Considering both token efficiency and reranking performance, we finally select Form 4 as our feature-based representation.
 However, in the second set of experiments where the number of documents is held constant, we also observe that feature-based representations may underperform full text.
-After all, offline information extraction may omit subtle but important cues present in the full text.
-Therefore, these results motivate a hybrid reranking strategy: using semantic features in the early stage to cover a broader range of candidates, followed by a refinement stage that reranks the top candidates using full-text inputs for more in-depth comparison.
-3 Methodology
-Based on the preliminary analysis, we propose a hybrid reranking framework in scientific domain.
-It proceeds in three steps: (i) offline extraction of structured semantic features; (ii) coarse-grained reranking using the features to select a subset of top candidates; and (iii) fine-grained reranking over the subset with full text inputs.
-3.1 Compact and Informative Representation
+After all, offline information extraction is not perfect.
+The process may omit subtle but important cues present in the full text.
+Therefore, these results motivate a balanced reranking strategy: using semantic features in the early stage to cover a broader range of candidates, followed by a refinement stage that reranks the top candidates using full text inputs for more in-depth comparison.
+4. Methodology
+Based on the preliminary analysis, we propose a new three-stage reranking framework in the scientific domain: (i) offline extraction of structured semantic features; (ii) coarse-grained reranking using the features to select a subset of top candidates; and (iii) fine-grained reranking over the subset with full text inputs.
+4.1. Compact and Informative Representation
 To construct compact, feature-based representations for each scientific document, we first perform document-level information extraction offline.
-As described in Section 2.3, we use zero-shot information extraction to obtain target semantic features: category, section, and keyword.
+As described in Section 3.3, we use LLMs to obtain target semantic features: category, section, and keyword.
 Formally, for a given document $p_{i}$, we obtain:
@@
 These features are designed to capture the core semantics of scientific documents in a token-efficient manner.
-The extraction is performed offline, and the results are cached and reused at inference time to avoid any additional runtime delay.
-The prompt templates for extracting different features can be found in Appendix A.
-3.2 Coarse Reranking w. Compact Features
+The extraction is performed offline, and the results are cached and reused during inference time to avoid any additional runtime delay.
+4.2. Coarse-grained Reranking w. Compact Features
 Given a query $q$, we obtain a document relevance ranking $\mathcal{C}=[p_{1},\,p_{2},\,\dots,\,p_{m}]$ from a first-stage retriever.
@@
 Specifically, for each unit (single section or keyword) $u$, we compute the cosine similarity with text embedding and select the top-5 keywords and the most relevant section.
-(Analysis on the number of keyword used can be found in Section 4.4) The final feature-based representation $r_{i}$ is formed as:
+(Analysis on the number of keyword used can be found in Section 2.) The final feature-based representation $r_{i}$ is formed as:
 $r_{i}=\text{Concat}\left[\text{Category}_{i},\,\text{Section}_{i}^{*},\,\text{Keywords}_{i}^{*}\right]$
@@
 With the feature-based representations $\mathcal{R}=[r_{1},\,r_{2},\,\dots,\,r_{m}]$, we then use the LLM to perform listwise reranking, producing a permutation $\sigma_{\text{feat}}=\text{LLM}(q,\,\mathcal{R})\in S_{m}$.
-Applying this permutation to the original candidate list $\mathcal{C}=[p_{1},\,p_{2},\,\dots,\,p_{m}]$ yields the reranked output for the coarse reranking:
+Applying this permutation to the original candidate list $\mathcal{C}=[p_{1},\,p_{2},\,\dots,\,p_{m}]$ yields the reranked output for the coarse-grained reranking:
 $\mathcal{C}_{\text{feat}}=[p_{\sigma_{\text{feat}}(1)},\,p_{\sigma_{\text{feat}}(2)},\,\dots,\,p_{\sigma_{\text{feat}}(m)}]$
@@
 $\mathcal{C}_{\text{seed}}=[p_{\sigma_{\text{feat}}(1)},\,\dots,\,p_{\sigma_{\text{feat}}(k)}],\quad k<m$
-Coarse reranking with compact document representations greatly expands the number of candidates that can be considered in the same LLM input.
-This broader coverage helps recover scientific documents that were initially assigned low scores by the first-stage retriever.
-3.3 Fine-grained Reranking w. Full Text
-In the second reranking stage, we refine the ranking over the seed set of candidates using full documents.
-Specifically, we take the seed set $\mathcal{C}_{\text{seed}}=[p_{1}^{\prime},\,p_{2}^{\prime},\,\dots,\,p_{k}^{\prime}]$, (we use to distinguish them from the first-stage input) obtained from the previous stage, and replace each compact representation with its original document text.
+Coarse-grained reranking with compact document representations greatly expands the number of candidates that can be considered in the same LLM input.
+This broader coverage helps recover scientific documents that were assigned low scores by the first-stage retriever.
+4.3. Fine-grained Reranking w. Full Text
+In the second reranking stage, we refine the ranking over the seed set of candidates using full text representation.
+Specifically, we take the seed set $\mathcal{C}_{\text{seed}}=[p_{1}^{\prime},\,p_{2}^{\prime},\,\dots,\,p_{k}^{\prime}]$, (we use to distinguish them from the first-stage input) obtained from the previous stage, and replace each compact representation with its original document.
 Let $\mathcal{T}=[t_{1},\,t_{2},\,\dots,\,t_{k}]$ denote the full text of the selected documents, where $t_{i}$ corresponds to the full text content of passage $p_{i}^{\prime}$.
@@
 $\mathcal{C}_{\text{final}}=[p_{\sigma_{\text{text}}(1)}^{\prime},\,\dots,\,p_{\sigma_{\text{text}}(k)}^{\prime}]$
-Fine-grained reranking with full-text inputs recovers details that may be lost during the information extraction process.
-Since the candidate set has already been narrowed down, full documents can now be used without exceeding the LLM’s context limit.
-Overall, this hybrid strategy effectively addresses the challenges of LLM-based listwise reranking in the scientific domain.
+Fine-grained reranking with full text inputs recovers details that may be lost during the information extraction process.
+Since the candidate set has already been narrowed down, full text representation can now be used without exceeding the LLM’s context limit.
+Overall, our strategy effectively addresses the challenges of LLM-based listwise reranking in the scientific domain.
 The coarse reranking stage expands the reranking pool, improving robustness to suboptimal first-stage retrieval in the scientific domain.
 The fine-grained reranking stage, in turn, preserves sufficient detail for precise relevance comparisons.
-Model Strategy LitSearch CSFCube
-N@5 N@10 M@5 M@10 R@5 R@10 N@5 N@10 M@5 M@10 R@5 R@10
-Initial Retriever
-Contriever Dense Retrieval 12.6 14.2 11.0 11.7 16.4 21.2 22.0 23.6 5.8 8.6 9.5 16.4
-Supervised, Model-Specific Methods
-[…]
-[…]
-[…]
-Zero-Shot, Model-Agnostic Reranking Strategies
+Table 1.
+Reranking performance of different reranking methods on LitSearch (Ajith et al., 2024), CSFCube (Mysore et al., 2021), NFCorpus (Boteva et al., 2016), SciFact (Wadden et al., 2020) and Trec-Covid (Voorhees et al., 2021) with nDCG@10 (N@10) and Recall@10 (R@10) as metrics.
+Higher scores under the same setting are bold.
+All values are shown as percentages.
+Reranking Model Reranking Strategy LitSearch CSFCube NFCorpus SciFact Trec-Covid [^1] Low because each query in Trec-Covid has far more than 10 relevant documents.
+N@10 R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10 R@10
+None First-stage retrieval only 14.2 21.2 23.6 16.4 32.8 15.5 67.7 81.3 59.6 1.6
 […]
@@
 […]
-Table 1: Reranking performance on the LitSearch and CSFCube reflected by nDCG (N@k), MAP (M@k) and Recall (R@k).
-We compare CoRank with supervised reranking models and zero-shot, model-agnostic strategy.
-4 Experiments
-In this section, we first outline the experimental setup, then present the reranking results, followed by ablation studies and further analysis.
-Detailed qualitative studies can be found in Appendix B.
-4.1 Experimental Setup
+[…]
+[…]
+[…]
+5. Experiments
+In this section, we first outline the experimental setup, then present the main reranking performance results, followed by ablation studies and further analysis.
+5.1. Experimental Setup
 Datasets & Metric.
-We evaluate different reranking methods on two high-quality scientific retrieval benchmarks: LitSearch (Ajith et al., 2024) and CSFCube (Mysore et al., 2021).
-The former is a benchmark of expert-annotated complex literature queries targeting recent ML and NLP papers.
-The latter is a human-annotated testbed for faceted query-by-example retrieval.
-Since reranking focuses on the top results, we follow standard practice (Sun et al., 2023a; Ma et al., 2023; Gangi Reddy et al., 2024; Liu et al., 2024c), reporting metrics for @5 and @10.
-We include nDCG, MAP, and Recall as our evaluation metrics.
-Models & Hyperparameters.
-For the reranking backbones, we use Qwen3-32B-Instruct (Qwen et al., 2024) as the representative open-source model, Gemini 2.0 Flash (Google DeepMind, 2025) and GPT-4.1-mini (OpenAI, 2024) as proprietary examples.
+We evaluate different reranking methods on five high-quality scientific retrieval benchmarks: LitSearch (Ajith et al., 2024), CSFCube (Mysore et al., 2021), NFCorpus (Boteva et al., 2016), SciFact (Wadden et al., 2020) and Trec-Covid (Voorhees et al., 2021), which collectively cover diverse scientific domains including computer science, biomedical and clinical information retrieval.
+Follow standard practice (Sun et al., 2023b; Ma et al., 2023; Gangi Reddy et al., 2024; Liu et al., 2024b), we report nDCG and Recall as our evaluation metrics for top-10 document.
+Models.
+For the reranking backbones, we use Qwen3-32B (Qwen et al., 2024) as the representative open-source model, Gemini 2.0 Flash (Google DeepMind, 2025) and GPT-4.1-mini (OpenAI, 2024) as proprietary examples.
 For generative parameters, we use a temperature of 1.0 and a fixed random seed of 42 for reproducibility.
-The API costs are detailed in Appendix C.
-For semantic feature extraction, we find that small open-source models are sufficiently effective; we therefore adopt Qwen3-8B-Instruct (Qwen et al., 2024) for all extraction tasks (Templates in Appendix A).
-For both first-stage retrieval and semantic similarity scoring for filtering features, we use Contriever (Izacard et al., 2021) following Gangi Reddy et al.
-(2024).
+For semantic feature extraction, we find that small open-source models are sufficiently effective; we therefore adopt Qwen3-8B (Qwen et al., 2024) for document feature extraction.
+To further assess the quality of information extraction outcome, we also conducted a human evaluation process shown in Section 5.3.5.
+For both first-stage retrieval and semantic similarity scoring for features filtering, we use Contriever (Izacard et al., 2021) following previous works (Reddy et al., 2023; Gangi Reddy et al., 2024).
+To ensure the robustness against different first-stage retrievers, We also evaluate the performance of our method with diverse retrievers in Section 5.3.3.
 Baselines.
-For supervised, model-specific baselines, we compare against RankVicuna (Pradeep et al., 2023a), RankZephyr (Pradeep et al., 2023b), and RankMistral (Liu et al., 2024c).
-These models are trained on large-scale general-domain datasets such as MS MARCO (Bajaj et al., 2016) and RankGPT (Sun et al., 2023a) to enhance reranking performance.
-For model-specific baselines, we use vanilla listwise reranking (Sun et al., 2023a) with zero-shot instruction-following.
-We also consider the sliding window strategy (Sun et al., 2023a; Ma et al., 2023), a test-time scaling (Xia et al., 2025) technique designed to enhance reranking performance.
-We report results both with and without the use of sliding windows.
-Although CoRank is orthogonal to this technique, we include a comparison in Section 4.3 to further show our effectiveness.
+Our method is a zero-shot, model-agnostic approach.
+Therefore, we primarily compare against prior reranking baselines under the same setting, most notably vanilla listwise reranking (Sun et al., 2023b) with zero-shot instruction-following.
+We also consider the sliding window strategy (Sun et al., 2023b; Ma et al., 2023), a test-time scaling technique (Xia et al., 2025) designed to expand the reranking context.
+Results are reported both with and without the sliding window strategy.
+We conduct a focused comparison between our method and the sliding window strategy in Section 5.3.2.
+We also include some reranking methods from the literature that represent the-state-of-the-art supervised LLM rerankers like RankVicuna (Pradeep et al., 2023a), RankZephyr (Pradeep et al., 2023b), and RankMistral (Liu et al., 2024b), which are trained on large-scale general-domain datasets such as MS MARCO (Bajaj et al., 2016).
+While these models differ from ours in both model size and supervision conditions, we report them for reference to illustrate the overall performance distribution.
+For a fairer comparison with supervised baselines, we evaluate our method using smaller-parameter LLMs (7B–8B) in Section 5.3.4, allowing us to isolate the effect of model size from supervision.
 Reranking Parameters.
-We follow the standardized setup (Sun et al., 2023a; Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) from prior work to unify reranking parameters.
-For vanilla reranking without the sliding window strategy, we include 20 full-text documents within the context.
-An exception is RankMistral (Liu et al., 2024c), which benefits from long-context training and is able to process 100 full-text documents in a single input.
-When using the sliding window strategy, we rank a total of 100 full-text documents by applying a window size of 20 with a step size of 10.
-For CoRank, we include 200 compact representations in the coarse reranking stage, followed by 20 full-text documents in the fine-grained reranking stage.
-4.2 Main Results
-Reranking Performance.
+We follow the standardized setup (Sun et al., 2023b; Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) from prior work to unify reranking parameters.
+For vanilla reranking, we include 20 full text documents within the context.
+An exception is RankMistral (Liu et al., 2024b), which benefits from long-context training and is able to process 100 full text documents in a single input.
+When using the sliding window strategy, we rank a total of 100 full text documents by applying a window size of 20 with a step size of 10.
+For CoRank, we include 200 compact representations in the coarse reranking stage, followed by 20 full text documents in the fine-grained reranking stage.
+5.2. Main Results
+5.2.1. Reranking Performance.
 We begin by evaluating the performance of different reranking methods on the target academic retrieval benchmarks, as shown in Table 1.
 First, we observe that all reranking methods outperform the first-stage dense retriever baseline, confirming the critical role of reranking in scientific document retrieval.
-However, while model-specific rerankers such as RankVicuna (Pradeep et al., 2023a) and zero-shot listwise baselines do yield noticeable improvements, their gains are relatively modest compared to results in general-domain settings (Sun et al., 2023a), particularly when the sliding window strategy is not used.
-In contrast, our proposed method, CoRank, is fully zero-shot, model-agnostic, and training-free, yet achieves strong and consistent gains across different LLM backbones.
-Notably, without using sliding windows, CoRank improves average nDCG@10 from 27.3 to 38.8, a relative improvement of over 40%.
-Even when combined with sliding window inputs, it still yields an average gain of +3.9 nDCG@10.
-We also find that both CoRank and the sliding window strategy independently lead to substantial performance gains (ours being larger and more efficient; see Section 4.3).
-This is likely because both methods expand the reranking scope, which is particularly valuable in scientific domains considering the suboptimal first-stage retrieval.
-These findings reinforce our earlier analysis of current limitations presented in Section 2.2.
-Ablation Studies.
+However, while supervised models such as RankVicuna (Pradeep et al., 2023a) do yield noticeable improvements, their gains remain limited, partly because these models are relatively small (7B–8B parameters) and thus less competitive despite having supervision.
+Meanwhile, current general LLMs already demonstrate strong instruction following capabilities for zero-shot reranking.
+Furthermore, our proposed method, CoRank, which is model-agnostic, and training-free, delivers even stronger and consistent gains across different LLM backbones.
+Without using sliding windows, CoRank improves the average nDCG@10 from 47.2 to 54.9, and when combined with sliding window inputs, it still provides an average gain of +2.0 nDCG@10 (from 54.1 to 56.2).
+We also find that both CoRank and the sliding window strategy independently lead to substantial performance gains (ours being larger and more efficient; see Section 5.3.2).
+This is likely because both methods expand the reranking scope, which is particularly valuable in the scientific domains considering the suboptimal first-stage retrieval.
+These findings also reinforce our earlier analysis of current limitations presented in Section 3.2.
+5.2.2. Ablation Studies.
 To evaluate the contribution of each component in our design, we conduct ablation studies on the LitSearch (Ajith et al., 2024) dataset using GPT-4.1-mini (OpenAI, 2024) as the backbone.
@@
 In addition, we ablate two key components of our pipeline: adaptive selection and fine-grained reranking, to understand their impact on overall performance.
-Model Strategy N@10 M@10 R@10
-Vanilla Full Text 26.1 25.2 27.3
-CoRank – Category 43.4 41.9 45.9
-CoRank – Section 45.4 43.8 48.2
-CoRank – Keywords 44.6 42.8 48.0
-CoRank – Selection 45.5 44.1 48.2
-CoRank – Fine. Rank. 41.1 38.7 46.7
-CoRank Full 46.3 44.7 49.4
-Table 2: Ablation results showing the impact of removing individual features and two key design components.
-As shown in Table 2, removing any single semantic component (category, section, or keywords) consistently degrades reranking performance, confirming that each feature contributes complementary information.
+Figure 4. Ablation study results. Foreground bars report performance after removing the labeled component; the translucent background bars indicate the full method score.
+[…]
+A horizontal bar chart shows ablation results on the LitSearch dataset (backbone: GPT-4.1-mini).
+The x-axis is “Score” from roughly 30 to 50.
+There are five rows labeled: w/o Keywords, w/o Section, w/o Category, w/o Fine.
+Rank.
+(removing fine-grained reranking), and w/o Selection (removing adaptive selection).
+For each row, two solid foreground bars report the metrics nDCG@10 (blue) and Recall@10 (red).
+Behind each, a faint background bar of the same color marks the full model (no ablation) score for reference.
+In every case the solid bars are shorter than the background bars, meaning performance drops when the component is removed; the largest drops are for removing Selection and Fine.
+Rank., while removing Category, Section, or Keywords yields smaller but consistent declines.
+As shown in Figure 4, removing any single semantic component (category, section, or keywords) consistently degrades reranking performance, confirming that each feature contributes complementary information.
 Category captures coarse-grained background context that tends to align more easily with the query but offers less specific detail.
 In contrast, section and keywords reflect finer-grained semantics that are harder to match but, when relevant, provide highly targeted signals.
-This balance underscores the value of combining semantic features across different levels of abstraction.
+This balance shows the value of combining semantic features across different levels of abstraction.
 Additionally, we find that both adaptive selection and fine-grained reranking contribute to the final performance.
 Adaptive selection ensures that the most query-relevant features are retained, while the fine-grained reranking stage substantially compensates for potential information loss in compact representations by recovering more nuanced evidence from the full text.
-4.3 Comparison with Sliding Window
-Sliding window strategy (Sun et al., 2023a) is a widely used (Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) method for overcoming the context length limitations of LLMs in listwise reranking.
+5.3. Further Analysis
+5.3.1. Hyperparameter Study
+We investigate the effect of key hyperparameters in CoRank using the GPT-4.1-mini (OpenAI, 2024) reranker on the LitSearch (Ajith et al., 2024) dataset, reporting nDCG@10, Recall@10, token usage, and estimated API cost [^2] Based on the pricing listed on the official OpenAI API Platform website..
+Our analysis focuses on understanding the trade-off between cost and performance, as well as the robustness of CoRank across a wide hyperparameter range.
+Number of Keywords.
+By default, CoRank extracts 30 keywords from each document and selects the top 5 most relevant ones via cosine similarity in the embedding space, concatenating them into the document representation for coarse-grained reranking.
+We vary the number of concatenated keywords from 0 to 20 and report the immediate results from coarse-grained reranking (Table 2).
+Table 2. Effect of the number of selected keywords on coarse-grained reranking: performance, token usage and cost.
+# Keyword N@10 R@10
+Token
+Usage
+Cost
+0 keyword 38.9 44.2 1.80M \$0.72
+1 keyword 40.4 46.3 1.95M \$0.78
+3 keywords 40.1 44.9 2.26M \$0.90
+5 keywords 41.1 46.7 2.57M \$1.03
+10 keywords 42.4 46.5 3.33M \$1.33
+15 keywords 42.6 48.6 4.10M \$1.64
+20 keywords 42.4 47.4 4.87M \$1.95
+Results show a clear cost–performance trade-off: adding more keywords linearly increases token usage while steadily improving coarse-grained reranking quality.
+However, performance gains diminish beyond 10 keywords, as later keywords contribute weaker relevance signals.
+Importantly, CoRank achieves strong performance even with substantially fewer keywords, indicating that it is robust to this hyperparameter and does not require precise tuning to remain competitive.
+Fine-Grained Pool Size.
+In the main experiments, the fine-grained reranker processes the top 20 candidates from the coarse stage to align with other baselines.
+Here, we vary the candidate pool from 5 to 100 documents and report the reranking performance of the final fine-grained stage (Table 3).
+Table 3. Effect of the fine-grained candidate pool size: performance, token usage and cost.
+Pool Size N@10 R@10
+Token
+Usage
+Cost
+5 docs 43.0 46.7 0.25M \$0.10
+10 docs 44.3 46.7 0.50M \$0.20
+20 docs 46.3 49.4 1.01M \$0.40
+40 docs 45.4 48.0 2.01M \$0.81
+60 docs 45.0 48.3 3.02M \$1.21
+80 docs 45.3 48.0 4.03M \$1.61
+100 docs 46.0 49.1 5.03M \$2.01
+We again observe a robust cost–performance trade-off: increasing the pool size up to 20 yields the most significant improvements, after which performance plateaus while token usage continues to grow.
+This plateau suggests that CoRank maintains stable effectiveness across a broad range of pool sizes, allowing practitioners to adjust this parameter flexibly according to budget without risking notable performance loss.
+Summary.
+These studies confirm that CoRank offers tunable hyperparameters that enable smooth navigation of cost–performance trade-offs, while maintaining consistently strong results.
+Its effectiveness is not narrowly dependent on exact hyperparameter values, alleviating concerns of fragility or heavy manual tuning.
+For example, varying the number of keywords between 5 and 20 results in only a 3.1% variance in nDCG@10, underscoring the robustness of the method.
+5.3.2. Comparison with Sliding Window
+Sliding window strategy (Sun et al., 2023b) is a widely used (Ma et al., 2023; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024) method for overcoming the context length limitations of LLMs in listwise reranking.
 Instead of ranking all candidates in a single prompt, it partitions the candidate list into overlapping windows and reranks each window independently from the bottom up.
-This method has been shown to notably improve reranking performance (Sun et al., 2023a).
+This method has been shown to notably improve reranking performance (Sun et al., 2023b).
 As noted in Section 5.1, we evaluate both the standalone and combined use of sliding windows with our method.
 While CoRank and sliding window are orthogonal and compatible, isolating them in comparison allows us to clearly demonstrate the superior efficiency and effectiveness of our approach when used alone.
-Method N@10 M@10 R@10 Token Usage Cost
-Vanilla 26.0 25.0 27.3 1.01M \$0.40
-Sliding 40.8 39.1 44.2 9.06M \$3.62
-CoRank 42.2 40.6 45.2 3.60M \$1.44
-Both 43.8 41.8 48.1 11.65M \$4.66
-Table 3: Comparison of performance, token usage, and API cost between CoRank and Sliding Window.
+Table 4. Comparison of performance, token usage, and API cost between CoRank and Sliding Window.
+Method N@10 R@10
+Token Usage
+Cost
+Vanilla 26.0 27.3 1.01M \$0.40
+Sliding 40.8 44.2 9.06M \$3.62
+CoRank 42.2 45.2 3.60M \$1.44
+Both 43.8 48.1 11.65M \$4.66
 We evaluate CoRank and the sliding window strategy on LitSearch and CSFCube on three tested models, reporting nDCG@10, total token consumption, and API cost for GPT-4.1-mini.
@@
 Moreover, combining these two methods can further boost performance, offering an advanced solution for complex retrieval scenarios.
-4.4 Hyperparameter Study
-We also study the effect of key hyperparameters in CoRank, using the GPT-4.1-mini (OpenAI, 2024) reranker on the LitSearch (Ajith et al., 2024) dataset.
-We report results in terms of nDCG@10, MAP@10, Recall@10, token usage, and estimated API cost.
-Number of Keywords.
-With its default configuration, CoRank extracts 30 keywords from each document and selects the top 5 most relevant ones based on cosine similarity from text embedding.
-These selected keywords are then concatenated into the document representation.
-Here, we evaluate the impact of the keyword selection on coarse reranking performance by varying the number of concatenated keywords: 0, 1, 3, 5, 10, 15, and 20.
-The results are in Table 4.
-# Keyword N@10 M@10 R@10 Token Usage Cost
-0 keyword 38.9 36.5 44.2 1.80M \$0.72
-1 keyword 40.4 38.0 46.3 1.95M \$0.78
-3 keywords 40.1 37.9 44.9 2.26M \$0.90
-5 keywords 41.1 38.7 46.7 2.57M \$1.03
-10 keywords 42.4 40.4 46.5 3.33M \$1.33
-15 keywords 42.6 40.1 48.6 4.10M \$1.64
-20 keywords 42.4 40.2 47.4 4.87M \$1.95
-Table 4: Effect of the number of selected keywords on coarse reranking: performance, token usage and cost.
-Our experiments reveal a trade-off between performance and cost when varying the number of keywords.
-In general, increasing the number of selected keywords leads to a linear increase in token cost, while also improving coarse reranking performance.
-However, we observe diminishing returns beyond 10 keywords.
-This is expected, as keywords are included in order of embedding similarity to the query—those added later tend to be less relevant and contribute weaker relevance signals.
-Fine-Grained Pool Size.
-In the main experiments, for fine-grained reranking, we select the top 20 candidates from the coarse reranking stage to align with other reranking baselines.
-Here, we explore the impact of varying this candidate selection size.
-Specifically, we evaluate performance when selecting 5, 10, 20, 40, 60, 80, and 100 candidates for fine-grained reranking.
-The results are shown in Table 5.
-Pool Size N@10 M@10 R@10 Token Usage Cost
-5 docs 43.0 41.3 46.7 0.25M \$0.10
-10 docs 44.3 42.8 46.7 0.50M \$0.20
-20 docs 46.3 44.7 49.4 1.01M \$0.40
-40 docs 45.4 43.9 48.0 2.01M \$0.81
-60 docs 45.0 43.4 48.3 3.02M \$1.21
-80 docs 45.3 43.8 48.0 4.03M \$1.61
-100 docs 46.0 44.4 49.1 5.03M \$2.01
-Table 5: Effect of the fine-grained candidate pool size: performance, token usage and cost.
-Experimental results show that increasing the fine-grained reranking candidate pool from 0 to 20 documents yields a favorable trade-off between token cost and performance.
-However, performance saturates beyond 20 candidates and remains stable up to 100 documents, despite the growth of token consumption.
-In theory, if an LLM’s long-context capability were fully effective, performance should continue to improve as more candidates are included.
-In practice, however, the observed plateau suggests that the effective context length is often much shorter than the theoretical context limit, which is also mentioned in previous studies (Hsieh et al., 2024; Kuratov et al., 2024).
-5 Related Work
-Classical Document Reranking.
-Document Reranking (Karpukhin et al., 2020; Ren et al., 2021, 2023) is a critical component in information retrieval (IR) (Baeza-Yates et al., 1999; Singhal et al., 2001), especially when high retrieval precision is required and the first-stage retriever alone is insufficient.
-Early reranking methods were typically lexical and probabilistic (Salton et al., 1975; Robertson and Jones, 1976), relying on term overlap between the query and document to adjust relevance scores.
-However, these approaches were limited by reliance on exact matching and failed to capture in-depth semantics (Karpukhin et al., 2020; Gao et al., 2021a).
-With advances in deep learning, neural reranking methods (Guo et al., 2019; Trabelsi et al., 2021), particularly those based on learning-to-rank (Cao et al., 2007; Liu, 2010) frameworks became prevalent.
-These models overcome previous limitations, capturing semantic similarity beyond surface-level term matches.
-The emergence of pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Reimers and Gurevych, 2019) has further transformed the field of document reranking (Nogueira et al., 2019a; Yates et al., 2021).
-Cross-encoder-based rerankers (Litschko et al., 2022), in particular, have demonstrated strong performance by modeling query-document relevance scores and optimizing pointwise, pairwise, or listwise loss functions (Zhuang et al., 2023).
-LLM-Based Listwise Document Reranking.
-Recently, large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) have revolutionized the field of NLP.
-With strong general knowledge and instruction-following (Ouyang et al., 2022) capabilities, LLMs offer a new paradigm for listwise document reranking.
-A classic proposal (Sun et al., 2023a) involves placing the query and a list of candidate documents into the context window and prompting the LLM to generate a reordered ranking.
-This zero-shot strategy has been shown to outperform traditional reranking baselines (Sun et al., 2023a; Ma et al., 2023).
-After that, LLM listwise reranking has been extended in multiple directions.
-Works like RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b) represent early explorations of open-source reranking LLMs.
-For inference time efficiency, FIRST (Gangi Reddy et al., 2024) modify listwise reranking by using only the first token decoding.
-RankMistral (Liu et al., 2024c) adopts long-context training to help LLMs adapt to larger document lists in a single input.
-Test-time scaling (Xia et al., 2025) also emerged as a common enhancement.
-For example, the sliding window strategy (Sun et al., 2023a; Ma et al., 2023) partitions the candidate list into overlapping chunks, enabling broader coverage within limited context.
-Building on this, Tang et al.
-(2024) enhance reranking quality through permutation consistency.
-ScaLR (Ren et al., 2024) further designs self-calibration to improve consistency across windows.
-Notably, two prior studies (Liu et al., 2024b; Li et al., 2024) also identified limitations of full-document representations, but their analyses are on general domain.
-Moreover, the alternative representations they propose are fundamentally different from ours—one (Liu et al., 2024b) adopts non-natural language embeddings, and the other (Li et al., 2024) relies on document chunking.
-To the best of our knowledge, our work is the first to highlight the unique limitations of full-document representations in scientific retrieval.
-We are also the first to leverage IE-based features as document representations for reranking.
-This makes our contribution novel and significant within the reranking literature.
-6 Conclusion
+Table 5.
+Human evaluation results for zero-shot LLM-based information extraction.
+The outcome achieves high accuracy, faithfulness, and completeness, with competitive preference rates compared to human annotations.
+Metric Definition Score
+Category Accuracy Proportion of LLM-assigned categories matching human-annotated ground truth 28/29 (96.6%).
+Topic Accuracy Proportion of LLM-assigned topics matching human-annotated ground truth 28/29 (96.6%).
+Section Faithfulness 1 - Proportion of generated sections that do not actually appear in the document. 192/197 (97.5%)
+Section Completeness 1 - Proportion of true sections in the document that the LLM failed to identify. 177/197 (89.8%)
+Keyword Faithfulness 1 - Proportion of generated keywords that do not actually appear in the document. 585/623 (93.9%)
+Keyword Completeness 1 - Proportion of true keywords in the document that the LLM failed to identify. 617/623 (99.0%)
+Win Rate Fraction of times annotators preferred the LLM’s annotation over the human’s. 7/15 (46.7%)
+5.3.3. Impact of First-Stage Retrievers
+Following previous works (Reddy et al., 2023; Gangi Reddy et al., 2024), we choose Contriever (Izacard et al., 2021) as our first-stage retriever in our main experiments.
+To further evaluate robustness to the choice of the first-stage retriever.
+We here additionally evaluate the performance of our method with BM25 (Robertson et al., 2009), Specter2 (Singh et al., 2023), GTR-T5-large (Ni et al., 2022) as first-stage retrievers.
+We run evaluation on LitSearch (Ajith et al., 2024) with Gemini 2.0 Flash (Google DeepMind, 2025).
+The results are shown in Figure 5.
+Figure 5. Comparison of reranking methods (None, Vanilla, CoRank) across four first-stage retrievers on LitSearch with Gemini 2.0 Flash.
+[…]
+The figure reports N@10 on LitSearch (Gemini 2.0 Flash) for four first-stage retrievers—Contriever, BM25, Specter2, and GTR-T5-large—under three reranking settings: None, Vanilla, and CoRank.
+For Contriever the scores are 14.2 (None), 31.4 (Vanilla), and 40.7 (CoRank); for BM25 they are 29.6, 49.1, and 52.8; for Specter2 they are 28.2, 47.2, and 53.7; and for GTR-T5-large they are 31.4, 48.6, and 49.5.
+CoRank is the best method for every retriever, improving over the raw first-stage (“None”) by +26.5, +23.2, +25.5, and +18.1 points respectively, and over the Vanilla reranker by +9.3, +3.7, +6.5, and +0.9.
+Under CoRank the ordering is Specter2 (53.7) ¿ BM25 (52.8) ¿ GTR-T5-large (49.5) ¿ Contriever (40.7), indicating consistent gains across retrievers and robustness to the first-stage choice.
+The first-stage retrievers exhibited differences in their standalone retrieval quality.
+However, our method consistently improved retrieval quality over the vanilla reranking baseline, regardless of which initial retriever was used.
+In other words, our gains are not confined to weak initial retrieval: CoRank adds value on strong first-stage retrievers as well, indicating that our method is robust to the choice of first-stage retriever.
+5.3.4. Impact of Reranking Backbones
+To further validate the robustness of our reranking method to different reranking backbones, we conducted additional experiments on smaller language models.
+While our main experiments employed LLMs like Qwen3-32B (Qwen et al., 2024) and Gemini 2.0 Flash (Google DeepMind, 2025), it remains a question whether our method generalizes effectively to more compact language models in the 7B–8B parameter range.
+We test two smaller models, Mistral-7B (Jiang et al., 2023) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) as reranking backbones and the results are shown in Table 6.
+Table 6. Reranking Performance with different small language models as backbones on LitSearch.
+Reranking Method Model Backbone N@10 R@10
+None None 14.2 21.2
+Vanilla Mistral-7B 20.9 24.6
+CoRank Mistral-7B 24.8 28.4
+Vanilla Llama-3.1-8B 22.1 25.3
+CoRank Llama-3.1-8B 26.0 31.2
+Vanilla RankMistral 20.1 24.3
+Vanilla RankVicuna 26.5 32.4
+Vanilla RankZephyr 34.1 39.0
+Our experiments confirmed two main findings.
+First, CoRank remained effective on smaller language models, consistently outperforming initial retrieval and vanilla reranking baseline.
+Second, although it did not surpass the supervised models in absolute performance, this outcome was expected, since the supervised methods require extensive training resources.
+For example, RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b) use GPT-3.5-distilled training data (1,000 K MS MARCO instances) and significant computational resources (320 GPU-hours on NVIDIA RTX A6000s).
+CoRank holds key practical advantages: it requires no additional training data or computational resources, works seamlessly as a plug-and-play method across different models and domains, and ensures data privacy.
+These strengths underscore CoRank’s value despite its lower absolute scores compared to supervised models.
+5.3.5. Human Evaluation on Feature Quality
+To more directly assess the quality of our zero-shot LLM-based information extraction—beyond the indirect evidence from retrieval performance and case in Figure 2, we conducted two human evaluation studies: (i) Human Verification and (ii) Human–LLM Comparison.
+Human Verification.
+Three domain experts independently evaluated the correctness of LLM-extracted categories, topics, sections, and keywords for a held-out set of 29 documents.
+For categories and topics, we measured accuracy, defined as the proportion of LLM-assigned labels that matched the human-annotated ground truth.
+For sections and keywords, which can have multiple correct instances per document, we assessed two complementary aspects: (i) Faithfulness – defined as 1 – hallucination rate, where hallucination rate is the proportion of elements generated by the LLM that do not actually appear in the source document; (ii) Completeness – defined as 1 – missing rate, where missing rate is the proportion of ground-truth elements in the document that the LLM failed to identify.
+This evaluation covered 29 categories, 29 topics, 197 section instances, and 623 keyword instances.
+All assessments were performed using the full text of each document and without exposure to other annotators’ judgments.
+Human–LLM Comparison.
+To measure subjective preference, annotators were presented with unlabeled pairs of annotations—one produced by a human annotator and one by the LLM—for 15 randomly selected documents.
+The identity of each annotation source was concealed, and annotators indicated which version they preferred for overall quality and informativeness.
+The primary metric was win rate, defined as the fraction of instances in which the LLM’s output was preferred over the human’s.
+Findings.
+As exhibited in Table 5.3.5, the LLM achieved high accuracy for category and topic assignments, strong faithfulness, and high completeness for both sections and keywords.
+The win rate in the comparison study indicates that, even without explicit optimization, the LLM’s outputs are competitive with human annotations.
+Overall, these results demonstrate that our zero-shot IE pipeline produces high-quality structured information across multiple content dimensions, providing direct evidence that the extraction process is not a limiting factor for our retrieval improvements.
+6. Conclusion
 In this work, we first identify fundamental challenges of standard listwise reranking in scientific retrieval.
 On the one hand, the first-stage retriever often fails to rank truly relevant documents high due to the lack of domain-specific training data.
-On the other hand, existing LLM-based listwise reranking methods typically operate over full-document inputs, which are substantial in terms of token usage and therefore limited to a small number of candidates.
-As a result, when the first-stage retrieval is suboptimal for scientific retrieval, the performance of standard reranking method is massively restricted.
+On the other hand, existing LLM-based listwise reranking methods typically operate over full-document inputs, which are substantial in token usage and therefore limited to a small number of candidates.
+As a result, when the first-stage retrieval is suboptimal for scientific retrieval, the performance of standard reranking method is restricted.
 To address this, we propose CoRank, a training-free and model-agnostic reranking framework for scientific retrieval.
-CoRank employs compact semantic features for coarse reranking.
+CoRank employs compact semantic features for coarse-grained reranking.
 This allows the LLM to consider a broader range of candidates within its context window.
-A subsequent fine-grained reranking stage then refines the top results using full-text inputs.
-Experiments on two scientific benchmarks show that CoRank significantly improves reranking performance.
-Overall, we are the first to explore semantic features as compact document representations for LLM reranking in scientific retrieval.
-Our results highlight the value of semantic feature extraction as a pre-analysis step in LLM listwise reranking in the scientific domain.
-Limitations
-Feature Extraction Quality.
-In our offline preprocessing stage, we apply zero-shot information extraction using LLMs to obtain semantic features such as categories for each document.
-This is a relatively simple approach.
-While it has proven effective based on both quantitative results and case studies, there is still potential room for improvement.
-For instance, one could explore using more specialized models or introducing multi-turn feedback mechanisms to enhance the quality of extracted features.
-That said, since the primary focus of our work is to address the reranking bottleneck in scientific retrieval, we did not dedicate significant effort to modifiying the IE component, especially observing that the basic extraction already yielded strong performance.
-Applicability Across Domains.
-This work focuses exclusively on scientific retrieval.
-The motivation and assumptions behind our method are specifically tailored to the characteristics of this domain, such as long-tail terminology and the limitations of first-stage retrievers.
-We do not evaluate our method in general-domain settings.
-The potential effectiveness and limitations of our method outside the scientific domain remain untested and are left for future work.
-Ethics Statement
-We conduct our experiments on widely recognized and publicly available scientific retrieval datasets.
-Our proposed method and findings do not pose any foreseeable harm to individuals or groups.
-Overall, we do not anticipate any significant ethical concerns with this work.
-Appendix A Prompt Templates
-In the preprocessing stage, our method uses Qwen3-8B-Instruct (Qwen et al., 2024) to perform zero-shot information extraction (Wei et al., 2023; Sainz et al., 2024), instructing the model to extract document-level semantic features such as categories (Sun et al., 2023b; Zhang et al., 2024), sections (Zhou et al., 2023), and keywords (Rose et al., 2010; Lee et al., 2023).
-During inference, we provide a zero-shot, model-agnostic reranking framework, and evaluate its performance on Qwen3-32B-Instruct (Qwen et al., 2024), Gemini 2.0 Flash (Google DeepMind, 2025), and GPT-4.1-mini (OpenAI, 2024).
-Here, we include the prompts used for different types of zero-shot document-level information extraction and model-agnostic reranking.
-A.1 Extracting Categories
-Prompt Template
-Please analyze this document for its topic and categories:
-{document} Provide a comprehensive analysis that includes: 1.
-The broad category (coarse-grained) this document belongs to 2.
-The specific category (fine-grained) within that broad category 3.
-A concise, title-like description of the document’s topic Deliver the analysis in one concise paragraph.
-A.2 Extracting Sections
-Prompt Template
-Identify 3-8 logical sections that would effectively organize this document’s content:
-{document} Generate appropriate subtitle-style headings for each section that would help structure the document. Sections should be comprehensive and cover the full scope of the content.
-A.3 Extracting Keywords
-Prompt Template
-Extract a comprehensive list of at least 30 diverse keywords and concepts from this document:
-{document} Generate as many diverse, relevant keywords and concepts as possible. Include both specific terms and broader conceptual themes.
-A.4 Generating Pseudo Queries
-Prompt Template
-Generate 20 diverse search queries that users might enter to find this document:
-{document} Create different types of queries that cover various aspects of the document content. Queries should be diverse in wording, length, and specificity.
-A.5 Reranking
-Prompt Template
-You are an LLM reranker, an intelligent assistant that can rank passages based on their relevancy to the query.
-I will provide you with {num} passages (either represented by full text, previous user query, keywords or structured analysis), each indicated by a numerical identifier [].
-Rank the passages based on their relevance to the search query: {query}.
-[1]{passage} [n]{passage} Search Query: {query}.
-Rank the {num} passages above based on their relevance to the search query.
-All the passages should be in descending order of relevance.
-The output format should be [passage_id] > [passage_id] > …, (If the full list is very long, generate at least 10) e.g., [4] > [2] > … Only respond with the ranking results, do not say any word or explain.
-Appendix B Case Studies
-B.1 Distribution of Positives and Negatives
-While metrics like nDCG@10 quantify the overall reranking quality, they do not indicate which specific reranking judgement contribute to the improvements.
-To better understand where our performance gains and limitations come from, we perform a qualitative analysis of how the position of ground-truth documents changes in the ranked list with or without CoRank.
-We compare CoRank with the vanilla zero-shot listwise reranker on LitSearch (Ajith et al., 2024) with GPT-4.1-mini (OpenAI, 2024).
-Based on the rank position of the ground-truth documents and whether they fall within the top-10 (i.e., Recall@10), we obtain the distribution of positive cases and negative cases.
-Figure 4: Distribution of Recall@10 positive and negative in Vanilla listwise reranking and CoRank.
-Analysis on Shared Positives.
-As shown in figure 4, CoRank successfully covers the majority of positive cases (69 out of 71) of the vanilla reranker.
-This is expected, as the fine-grained reranking stage in CoRank inherits the full-text completeness in document representation of the vanilla listwise reranking.
-Analysis on Exclusive Positives.
-At the same time, CoRank exclusively attains 58 positive cases.
-Analyzing these cases, we find that most (57 out of 58) were missed by the vanilla reranker because their first-stage retrieval ranks were too low (i.e., outside the top 20).
-This supports our analysis in Section 2.2 on the key challenge in scientific-domain reranking: truly relevant documents may be excluded from reranking due to suboptimal first-stage retrieval results.
-CoRank mitigates this issue with the compactness of feature-based representations in the coarse reranking stage, allowing more candidates to be considered and increasing the likelihood of recovering relevant documents.
-Next, we present full instances of both positive and negative examples, along with their ranks in the first-stage retrieval, coarse reranking stage, and fine-grained reranking stage.
-This allows us to analyze the underlying causes of success or failure in each case.
-B.2 Positive Example
-This example clearly demonstrates the advantages of our method.
-The ground-truth document in this case involves long-tailed scientific concepts such as CGA and VAE, which are not well captured by the first-stage retriever.
-As a result, it is ranked only 92nd in the initial retrieval, making it inaccessible to vanilla rerankers that operate on the top 20 candidates.
-Even with the sliding window strategy (Sun et al., 2023a; Ma et al., 2023), the document would need to consistently win across 9 windows to enter the top-10, highlighting the limitations of existing reranking methods in scientific domains, as discussed earlier.
-In contrast, our method successfully ranks this document as the top candidate during the coarse reranking stage.
-This illustrates two key points: (1) our feature-based representation is highly compact, allowing us to expand the reranking pool to 200 candidates and include low-ranked but relevant documents such as this one; and (2) it is also sufficiently informative, enabling the reranker to recognize and promote the document directly to the top of the list once included.
-Query
-Can you recommend a foundational paper that provides a scalable framework for generating English sentences with controllable semantic and syntactic attributes for the purpose of augmenting datasets in NLP tasks?
-Ground Truth Document Control, Generate, Augment: A Scalable Framework for Multi-Attribute Text Generation: We introduce CGA, a conditional VAE architecture, to control, generate, and augment text.
-CGA is able to generate natural English sentences controlling multiple semantic and syntactic attributes by combining adversarial learning with a context-aware loss and a cyclical word dropout routine.
-We demonstrate the value of the individual model components in an ablation study.
-The scalability of our approach is ensured through a single discriminator, independently of the number of attributes.
-We show high quality, diversity and attribute control in the generated sentences through a series of automatic and human assessments.
-As the main application of our work, we test the potential of this new NLG model in a data augmentation scenario.
-In a downstream NLP task, the sentences generated by our CGA model show significant improvements over a strong baseline, and a classification performance often comparable to adding same amount of additional real data.
-First-Stage Retrieval Ranking 92 of 14256
-Compact Representation Natural Language Processing (NLP) -> Text Generation and Neural Machine Translation -> Conditional VAE-Based Framework for Controllable and Scalable Multi-Attribute Text Generation with Applications in Data Augmentation: CGA for Data Augmentation in NLP Tasks (Text Generation, Multi-Attribute Control, Data Augmentation, Semantic Attributes)
-Coarse Reranking 1 of 200
-Fine-grained Reranking 1 of 20
-B.3 Negative Example
-The example above illustrates a case where our method fails to rank the ground-truth document highly—not due to representation quality, but because the document was ranked extremely low (319th) by the first-stage retriever.
-At such a low position, it is not visible to vanilla reranking methods, nor to those with extended context windows such as the sliding window strategy (Sun et al., 2023a; Ma et al., 2023) or our Method.
-As a result, even though our compact representation accurately captures the document’s semantic content, it never enters the reranking process and thus has no opportunity to be promoted to a higher position.
-Query
-Are there any studies that explore post-hoc techniques for hallucination detection at both the token- and sentence-level in neural sequence generation tasks?
-Ground Truth Document Detecting Hallucinated Content in Conditional Neural Sequence Generation: Neural sequence models can generate highly fluent sentences, but recent studies have also shown that they are also prone to hallucinate additional content not supported by the input.
-These variety of fluent but wrong outputs are particularly problematic, as it will not be possible for users to tell they are being presented incorrect content.
-To detect these errors, we propose a task to predict whether each token in the output sequence is hallucinated (not contained in the input) and collect new manually annotated evaluation sets for this task.
-We also introduce a method for learning to detect hallucinations using pretrained language models fine tuned on synthetic data that includes automatically inserted hallucinations Experiments on machine translation (MT) and abstractive summarization demonstrate that our proposed approach consistently outperforms strong baselines on all benchmark datasets.
-We further demonstrate how to use the token-level hallucination labels to define a fine-grained loss over the target sequence in low-resource MT and achieve significant improvements over strong baseline methods.We also apply our method to word-level quality estimation for MT and show its effectiveness in both supervised and unsupervised settings 1.
-Compact Representation Natural Language Processing -> Hallucination Detection in Neural Sequence Generation -> Token-Level Hallucination Detection in Conditional Text Generation: detecting hallucination using token-level classification and pretrained language models (Hallucination Detection, Neural Sequence Generation, Token-Level Classification, Machine Translation (MT))
-First-Stage Retrieval Ranking 319 of 14256
-Coarse Reranking N/A
-Fine-grained Reranking N/A
-Appendix C Proprietary Model API Cost
-As part of our evaluation of different reranking methods, we tested two commercial LLMs: Gemini 2.0 Flash (Google DeepMind, 2025) and GPT-4.1-mini (OpenAI, 2024).
-Here, we report the API costs incurred in the main experiments when querying these models.
-The costs associated with experiments in the Further Analysis section have already been presented as part of the token efficiency comparison.
-Model Price Setting # Token Cost (\$)
-[…]
-[…]
-[…]
-[…]
-Total Cost 29.61M \$2.96
-[…]
-[…]
-[…]
-[…]
-Total Cost 29.61M \$11.84
-Table 6: Reranking token usage and cost across different settings for Gemini 2.0 Flash and GPT-4.1-mini (OpenAI, 2024). Prices are per 1 million tokens.
-As shown in Table 6, across all different settings in the main experiments, we spent \$2.96 on Gemini-2.0-Flash (Google DeepMind, 2025) and \$11.84 on GPT-4.1-mini (OpenAI, 2024), resulting in a total API cost of \$14.80.
-Appendix D Future Work
-In the previous section, we identified two main limitations of our current approach.
-These suggest promising directions for future research that could further improve the generality and effectiveness of our method.
-D.1 Improving Feature Extraction
-Our current pipeline uses zero-shot information extraction (Wei et al., 2023) with general-purpose LLMs (Qwen et al., 2024) to generate semantic features such as categories (Sun et al., 2023b; Zhang et al., 2024) and keywords (Rose et al., 2010; Lee et al., 2023).
-While this approach is simple and has proven effective, future work could explore more advanced techniques to enhance extraction quality.
-For example, using domain-specific LLMs (Taylor et al., 2022; Xie et al., 2023) or introducing multi-turn feedback (Shinn et al., 2023) may yield richer and more accurate representations, further improving downstream reranking performance.
-D.2 Extending to General Domains
-Our method is designed specifically for scientific retrieval (Lawrence et al., 1999; White et al., 2009), where challenges like long-tail terminology (Kang et al., 2024a) and limited first-stage recall (Kim et al., 2023) are particularly severe.
-An important direction for future work is to investigate the applicability of our framework in general-domain retrieval settings (Zhu et al., 2023; Ai et al., 2023).
-This includes evaluating how well the feature-based representation and the hybrid reranking strategy transfer beyond the scientific context, and identifying any necessary adaptations.
+A subsequent fine-grained reranking stage then refines the top results using full text inputs.
+Experiments on five scientific datasets show that CoRank significantly improves the reranking performance.
+Overall, we explore semantic features as compact document representations for LLM reranking in scientific retrieval.
+Our results demonstrate how integrating semantic feature extraction into the retrieval pipeline creates a synergistic interaction between information extraction and reranking, leading to more effective scientific literature retrieval.
+Ethical Impact Statement
+This work focuses on improving scientific document retrieval using publicly available datasets, and we do not foresee direct harm to individuals or groups.
+Nonetheless, LLM-based reranking methods can inherit biases from underlying corpora, potentially overrepresenting certain disciplines or communities.
+They could also, in real-world deployments, surface sensitive or dual-use technical content.
+To mitigate these risks, we recommend applying fairness audits, integrating content filters for high-risk topics, and maintaining human oversight in the interpretation of retrieved results.
+Additionally, caution should be taken to avoid overreliance on automated rankings, ensuring they complement rather than replace critical human judgment.